### v1

# User-Creator Feature Dynamics in Recommender Systems with Dual Influence

Abstract
Recommender systems present relevant contents to users and help content creators reach their target audience. The dual nature of these systems influences both users and creators: users’ preferences are affected by the items they are recommended, while creators are incentivized to alter their contents such that it is recommended more frequently. We define a model, called user-creator feature dynamics, to capture the dual influences of recommender systems. We prove that a recommender system with dual influence is guaranteed to polarize, causing diversity loss in the system. We then investigate, both theoretically and empirically, approaches for mitigating polarization and promoting diversity in recommender systems. Unexpectedly, we find that common diversity-promoting approaches do not work in the presence of dual influence, while relevancy-optimizing methods like top-$k$ recommendation can prevent polarization and improve diversity of the system.
Recommender Systems, Preference Dynamics, Dual Influence
1 Introduction
From restaurant selection, video watching, to apartment renting, recommender systems play a pivotal role across a plethora of real-world domains. These systems match users with items they like, and help creators (those producing the items) identify their target audiences. Nevertheless, behind such success, concerns have emerged regarding possible harmful outcomes of recommender systems, in particular, filter bubbles (Masrour et al., 2020; Aridor et al., 2020) and polarization (Santos et al., 2021) – outcomes with insufficient recommendation diversity and creation diversity. Recommendation diversity, meaning the diversity of the contents recommended to a user, is key to users’ engagement and retention on the platform. Meanwhile, creation diversity, meaning the variety of content created on the platform, is a determinant of the platform’s long-term health. In extreme cases, insufficient creation diversity can lead to consensus or polarization, where the latter can cause conflict and hatred, diminish people’s mutual understanding, and cause societal crises. Therefore, from both business and social responsibility perspectives, championing and improving diversity in recommender systems is equally important as optimizing recommendation relevance.
There is increasing emphasis in academia and industry on investigating and improving the diversity of recommender systems, combating filter bubbles and polarization. Popular diversity-boosting approaches include applying post-processing procedures such as re-ranking (Carbonell & Goldstein, 1998; Ziegler et al., 2005) and setting diversity-aware objectives in addition to relevance maximization (Su et al., 2013; Zhang & Hurley, 2008; Hurley, 2013; Wilhelm et al., 2018; Cheng et al., 2017). These methods aim to increase the recommendation diversity for users. Assuming that the contents on the platform are static, these methods have been shown to bring diversity gain to the system.
However, an important aspect that is overlooked in the aforementioned approaches is that: users and contents on a recommendation platform are not static entities – they can be influenced by the recommendation made by the system. In content creation platforms like YouTube, TikTok, and Twitter, recommendations naturally affect both content users and content creators. It is well known that the exposure to recommended items can shift a user’s preference (Jiang et al., 2019; Dean & Morgenstern, 2022). On the other hand, the creators have the incentive to change their creation styles constantly to attract their audience better (and to make more profits from the platform) (Eilat & Rosenfeld, 2023; Hron et al., 2023; Jagadeesan et al., 2024). While the effects of recommendation on either users or creators have been investigated separately, no previous work considers both effects to our knowledge. The dual influence of recommendation on users and creators causes complicated dynamics where users and creators interact and their preferences evolve together. Such evolution might exacerbate filter bubbles and polarization. Whether the aforementioned diversity-boosting approaches still work in a dynamic environment with dual influence is questionable.
The first contribution of our work is to define a novel, natural dynamics model that captures the dual influence of a recommender system on users and creators, which we call user-creator feature dynamics (Section 2). We leverage the users’ and items’/creators’ embedding vectors to represent their preferences and creation styles, and use cosine similarity to characterize the relevance of creations and users’ interests (which is common in the recommender system literature and practice). This model allows us to formally reason about the impact of various design choices on the diversity of a recommender system with dual influence.
Our second contribution is to demonstrate that, under realistic conditions, the user-creator feature dynamics of any recommender system with dual influence is guaranteed to converge to polarization (Section 3), i.e., the preferences of users and the contents of creators will be tightly clustered into two opposite groups (bringing the diversity of the system to nearly zero). We demonstrate that this phenomenon still occurs even after applying diversity-boosting approaches to the recommender system.
Then, (in Section 4) we investigate some real-world designs of recommendation algorithms in order to look for techniques that mitigate polarization. Interestingly, we find that some common efficiency-improving methods, such as top-$k$ recommendation, can both prevent the system from polarization and improve the creation diversity. We also provide empirical results on both synthetic and real-world (MovieLens) data that support our theoretical claims; see Section 5 for a summary. Section 6 concludes.
Additional Related Works
The most related works have been mentioned above. App. A offers additional discussions on three strands of literature: diversity in recommender systems, opinion dynamics, and performative prediction.
2 Model: User-Creator Feature Dynamics
We define a dynamics model for user preferences and content/creator features in a recommender system. Let $\bm{U}^{t}=[\bm{u}_{j}^{t}]_{j=1}^{m}=[\bm{u}_{1}^{t},\ldots,\bm{u}_{m}^{t}]\in{\mathbb{R}}^{d\times m}$ be a population of $m$ users and $\bm{V}^{t}=[\bm{v}_{i}^{t}]_{i=1}^{n}=[\bm{v}_{1}^{t},\ldots,\bm{v}_{n}^{t}]\in{\mathbb{R}}^{d\times n}$ be a population of $n$ creators at time $t$, where each vector $\bm{u}_{j}^{t},\bm{v}_{i}^{t}\in\mathbb{S}^{d-1}$ represent the preference/feature vector of each user and creator respectively, assumed to be on the unit sphere $\mathbb{S}^{d-1}$ with $\ell_{2}$-norm. Then $(\bm{U}^{t},\bm{V}^{t})$ denotes the state of the dynamics at time $t$. The dynamics evolve as follows at each time step $t\geq 0$:
1) Recommendation: Each user $j\in[m]$ is recommended a creator, where creator $i\in[n]$ is chosen with a probability
$p_{ij}^{t}=p_{ij}^{t}(\bm{U}^{t},\bm{V}^{t}).$ (1)
While we allow a wide array of different functions $p_{ij}^{t}(\cdot)$, a common example of such functions is the so-called softmax function:
$p_{ij}^{t}=\mathrm{softmax}(\bm{u}_{j}^{t},\bm{V}^{t};\beta)=\tfrac{\exp(\beta\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle)}{\sum_{i=1}^{n}\exp(\beta\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle)}.$ (2)
A larger $\beta$ means that the recommendation is more sensitive to the relevance of a creator to a user, measured by $\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle$.
2) User update: After recommendation, each user $j\in[m]$ updates their feature vector $\bm{u}_{j}^{t}$, based on which creator, say $i_{j}^{t}$, was recommended to them:
$\bm{u}_{j}^{t+1}=\mathcal{P}\big{(}\bm{u}_{j}^{t}+\eta_{u}f(\bm{v}_{i_{j}^{t}}^{t},\bm{u}_{j}^{t})\bm{v}_{i_{j}^{t}}^{t}\big{)}.$ (3)
Here, $\eta_{u}>0$ is a parameter controlling the rate of update, $f(\bm{v}_{i},\bm{u}_{j})$ is a function that quantifies the impact of creator $i$’s content on user $j$, and $\mathrm{\mathcal{P}}(\bm{x})=\frac{\bm{x}}{\|\bm{x}\|_{2}}$ is the projection onto the unit sphere. Our user update model generalizes (Dean & Morgenstern, 2022).
3) Creator update: Creators also update their feature vectors based on which users are recommended their content. For each creator $i\in[n]$, let $J_{i}^{t}=\{j:i_{j}^{t}=i\}$ be the set of users being recommended creator $i$, then $\bm{v}_{i}^{t}$ is updated by:
$\bm{v}_{i}^{t+1}=\mathcal{P}\Big{(}\bm{v}_{i}^{t}+\frac{\eta_{c}}{|J_{i}^{t}|}\sum_{j\in J_{i}^{t}}g(\bm{u}_{j}^{t},\bm{v}_{i}^{t})\bm{u}_{j}^{t}\Big{)},$ (4)
where $\eta_{c}>0$ is an update rate, and $g(\bm{u}_{j},\bm{v}_{i})$ is a function that quantifies the impact of user $j$ on creator $i$. Our creator update model generalizes (Eilat & Rosenfeld, 2023).
Impact functions $f$ and $g$
Our results apply to any impact functions $f$ and $g$ that satisfy the following natural assumptions. First, $f(\bm{v}_{i},\bm{u}_{j})$ and the inner product $\langle\bm{v}_{i},\bm{u}_{j}\rangle$ have the same sign: $f(\bm{v}_{i},\bm{u}_{j})\text{ is }{\tiny\begin{cases}>0&\text{ if }\langle\bm{v}_{i},\bm{u}_{j}\rangle>0\\ <0&\text{ if }\langle\bm{v}_{i},\bm{u}_{j}\rangle<0\\ =0&\text{ if }\langle\bm{v}_{i},\bm{u}_{j}\rangle=0.\end{cases}}$ This means that if a user likes the content ($\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}\rangle>0$), then the user vector $\bm{u}_{j}^{t}$ will be updated towards the direction of the creator vector $\bm{v}_{j}^{t}$. If the user dislikes the content ($\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}\rangle<0$), then the user vector $\bm{u}_{j}^{t}$ will move away from $\bm{v}_{j}^{t}$. Further assume upper and lower bounds on $|f|$:
$|f(\bm{v}_{i},\bm{u}_{j})|\leq 1,\quad~{}|f(\bm{v}_{i},\bm{u}_{j})|\geq L_{f}>0\text{ when $\langle\bm{v}_{i},\bm{u}_{j}\rangle\neq 0$}.$
The lower bound $|f(\bm{v}_{i},\bm{u}_{j})|\geq L_{f}$ means that the exposure to an item that a user likes or dislikes always has some non-negligible impact on the user’s preference. For example, $f(\bm{v}_{i},\bm{u}_{j})=\mathrm{sign}(\langle\bm{v}_{i},\bm{u}_{j}\rangle)a+b\langle\bm{v}_{i},\bm{u}_{j}\rangle$ satisfies both assumptions when $L_{f}=a>0$ and $b\geq 0$.
For $g$, likewise assume that its sign is the same as $\langle\bm{u}_{j},\bm{v}_{i}\rangle$: $g(\bm{u}_{j},\bm{v}_{i})\text{ is }{\tiny\begin{cases}>0&\text{ if }\langle\bm{u}_{j},\bm{v}_{i}\rangle>0\\ <0&\text{ if }\langle\bm{u}_{j},\bm{v}_{i}\rangle<0\\ =0&\text{ if }\langle\bm{u}_{j},\bm{v}_{i}\rangle=0.\end{cases}}$ Intuitively, this captures the incentive of a creator who tries to maximize the average ratings from users who are recommended their items. On video platforms for example, if the creators are rewarded based on the average rating of their videos, they will try to reinforce their creation style based on the users who give positive feedback ($\langle\bm{u}_{j},\bm{v}_{i}\rangle>0$) so that their creations are more likely to be recommended to those users. Meanwhile, they will also change their creation style based on negative feedback ($\langle\bm{u}_{j},\bm{v}_{i}\rangle<0$), but in the opposite direction of the negative-feedback users’ interests, so that their creations are less likely to be recommended to those users. A particular example of $g$ is the sign function $g(\bm{u}_{j},\bm{v}_{i})=\mathrm{sign}(\langle\bm{u}_{j},\bm{v}_{i}\rangle)\in\{-1,0,1\}$. We will focus on the sign function $g$ to simplify the theoretical presentation. All our other results hold for general $g$ functions.
3 Unavoidable Polarization
Having defined the user-creator feature dynamics in a recommender system with dual influence, we theoretically study how such dynamics evolve. Our main result is: if every creator can be recommended to every user with some non-zero probability, then the dynamics must eventually polarize.
Definition 3.1 (consensus and bi-polarization).
Let $R>0$. The dynamics $(\bm{U}^{t},\bm{V}^{t})$ is said to reach:
  • $R$-consensus if there exists a vector $\bm{c}\in{\mathbb{R}}^{d}$ such that every feature vector is $R$-close to $\bm{c}$: $\forall\bm{u}_{j}^{t},\|\bm{u}_{j}^{t}-\bm{c}\|_{2}\leq R$ and $\forall\bm{v}_{i}^{t},\|\bm{v}_{i}^{t}-\bm{c}\|_{2}\leq R$.
  • $R$-bi-polarization if there exists a vector $\bm{c}\in{\mathbb{R}}^{d}$ such that every feature vector is $R$-close to $+\bm{c}$ or $-\bm{c}$: $\forall\bm{u}_{j}^{t}$, $\|\bm{u}_{j}^{t}-\bm{c}\|_{2}\leq R$ or $\|\bm{u}_{j}^{t}+\bm{c}\|_{2}\leq R$ holds, and $\forall\bm{v}_{i}^{t}$, $\|\bm{v}_{i}^{t}-\bm{c}\|_{2}\leq R$ or $\|\bm{v}_{i}^{t}+\bm{c}\|_{2}\leq R$ holds.
Consensus is any state where all users and creators have similar feature vectors (up to difference $R$), implying that they have similar preferences. Bi-polarization is any state where all users and creators are clustered into two groups with exactly opposite features (e.g., Republicans vs Democrats).
Proposition 3.2.
Bi-polarization states are absorbing: namely, once the dynamics reaches $R$-bi-polarization with some $R\in[0,1]$, it will satisfy $R$-bi-polarization forever. The same holds for consensus.
A natural property of a recommender system is that every creator can be recommended to every user with some non-zero probability: $p_{ij}^{t}\geq p_{0}>0$. This is satisfied by the softmax function, which is a rough model of real-world recommendations (Covington et al., 2016): $p_{ij}^{t}=\tfrac{\exp(\beta\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle)}{\sum_{i=1}^{n}\exp(\beta\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle)}\geq\tfrac{\exp(-\beta)}{n\exp(\beta)}=p_{0}>0$. Moreover, many large-scale real-world recomendation systems (e.g., Yahoo! (Li et al., 2010) and Kuaishou (Gao et al., 2022)) intentionally insert small random traffic to improve recommendation diversity or to explore users’ interests (Judith Möller & van Es, 2018; Yang et al., 2018), which will cause all recommendation probabilities to be non-zero. We show that, however, a recommender system satisfying $p_{ij}^{t}\geq p_{0}>0$ must converge to polarization, under some additional conditions on the users’ and creators’ update rates:
Theorem 3.3.
Suppose $g(\bm{u}_{j},\bm{v}_{i})=\mathrm{sign}(\langle\bm{u}_{j},\bm{v}_{i}\rangle)$, the update rates $\eta_{u}<\frac{1}{2}$ and $\eta_{c}\leq\frac{\eta_{u}L_{f}}{2}$, and the recommendation probability $p_{ij}^{t}\geq p_{0}>0,\forall i,j,t$. Then, from almost all initial states, the dynamics $(\bm{U}^{t},\bm{V}^{t})$ will eventually reach $R$-consensus or $R$-bi-polarization for any $R>0$.
In other words, if the users’ and creators’ updates are not too fast and the recommendation probabilities are non-zero, then all users and creators will eventually converge to at most two clusters (regardless of the feature dimension $d$). Since creators in one cluster produce similar contents, users in such a polarized system can never receive diverse recommendations. This means that the naïve attempt of imposing $p_{ij}^{t}\geq p_{0}>0$ cannot improve the diversity of a recommender system with dual influence. The conditions on the update rates $\eta_{u},\eta_{c}$ are only assumed to simplify the proof of Theorem 3.3. Our experiments (in Appendix B) show that polarization still occurs even without those conditions.
4 Discussions on Real-World Designs
The result in the previous section raises a natural question of whether polarization can be mitigated through careful design of recommender systems. We discuss 4 types of designs: top-$k$ recommendation, truncation, diversity-boosting, and uniform traffic.
(1) Top-$k$ Recommendation
A prevalent practice in modern two-stage recommendation algorithms on large-scale platforms, such as YouTube (Covington et al., 2016), is to first filter out items that are unlikely to be relevant to a user, then make recommendations from the remaining items. This practice reduces the computation cost and improves the relevancy of recommendations. Interestingly, we show that such a practice can also improve the long-term diversity of a recommender system with dual influence.
In particular, we consider the top-$k$ recommendation policy where each user is recommended only the $k$ most relevant creators, so $p_{ij}^{t}=0$ if $i$ is not one of $k$ creators $i^{\prime}$ that maximize $\langle\bm{v}_{i^{\prime}}^{t},\bm{u}_{j}^{t}\rangle$. We show that top-$k$ recommendation allows the dynamics to potentially converge to states with multiple clusters, instead of bi-polarization.
Proposition 4.1.
With top-$k$ recommendation, there exist absorbing states that form $\lfloor n/k\rfloor$ clusters (i.e., once the system forms $\lfloor n/k\rfloor$ clusters, it forms $\lfloor n/k\rfloor$ clusters forever).
In contrast, Theorem 3.3 said that a recommender system where every creator can be recommended to every user ($p_{ij}^{t}>0$) is doomed to polarize. But with top-$k$ recommendation where some $p_{ij}^{t}=0$, polarization can be avoided. Experiments in Appendix B.3 also show that top-$k$ recommendation can reduce polarization and improve diversity.
(2) Truncation
Besides top-$k$ recommendation, truncation is another way to filter out irrelevant creators: set a threshold $\tau\in[-1,1]$ such that any user-creator pair with inner product $\langle\bm{u}_{i},\bm{v}_{j}\rangle<\tau$ is not recommended. Increasing $\tau$ is to similar to decreasing $k$ in top-$k$ recommendation, which improves recommendation relevance. Our experiments (in Appendix B) show that truncation at a large threshold helps to reduce polarization and improve diversity in a recommender system with dual influence.
(3) Diversity Boosting
Diversity boosting tries to explore users’ interests and improve users’ experience by diversifying recommendation. For example, when making recommendations, the model optimizes the objective $h_{rel}(\langle\bm{u}_{i},\bm{v}_{j}\rangle)+\rho\cdot h_{div}(list_{i},\bm{v}_{j}),$ where $h_{rel},h_{div}$ rewards the recommendation relevance and diversity respectively and $list_{i}$ records the recent list of recommended items to user $i$. $h_{div}$ can take a simple form of $\sum_{j^{\prime}\in list_{i}}1-\langle\bm{v}_{j^{\prime}},\bm{v}_{j}\rangle$, and $\rho>0$ controls the strength of diversity-boosting. Despite being successful with fixed items and user preferences, this design cannot prevent bi-polarization in our dual-influence dynamics, since the conditions in Theorem 3.3 are still satisfied. Experiments in App. B.2 support our claim.
(4) Uniform Traffic
Another method proposed in previous works (Judith Möller & van Es, 2018; Gao et al., 2022; Borgs et al., 2023) to improve the diversity of recommendation is to add a small fraction of uniform traffic to the personalized recommendation. As we argued in Sectin 3, this method could cause polarization and diversity loss in a recommender system with dual influence. Such an observation is striking as it demonstrates that optimizing for recommendation diversity in a static setting can ultimately lead to a huge loss of the system diversity in the long run.
5 Observations from Experiments
We conducted experiments on synthetic data and real-world data (MovieLens 20M (Harper & Konstan, 2015)). We simulated the user-creator feature dynamics for hundreds of time steps and observed the changes of four measures for the diversity of the system (Creator Diversity, Recommendation Diversity, Recommendation Relevance, and Tendency to Polarization) under different choices of parameters and real-world designs. Here are our three main observations:
  • Smaller sensitivity $\beta$ in the softmax function, larger creator update rates $\eta_{c}$, and larger user update rate $\eta_{u}$ cause faster polarization and worse diversity for the system. Notably, uniform recommendation ($\beta=0$), with the most diverse recommendation for the users in the short term, leads to severe polarization in the long run.
  • Top-$k$ recommendation with a small $k$ (Fig. 1) and truncation with a large threshold $\tau$ can alleviate polarization and improve the diversity of the system in the long run.
  • Diversity boosting methods, despite having a diversity-aware objective, fasten polarization and worsen the diversity of the system in the long run.
In summary, we found that, due to the dual influence of recommender systems, naïve diversity-promoting methods (like decreasing $\beta$ and using diversity-aware objective) cannot improve and even hurt the long-term diversity of the system; this is consistent with our Theorem 3.3 because those methods satisfy $p_{ij}^{t}\geq p_{0}>0$. Methods targeting efficiency and relevancy of recommendation (like top-$k$ recommendation), not satisfying $p_{ij}^{t}\geq p_{0}>0$, can instead improve the long-term diversity of the system.
Figure 1: Changes of diversity measures over time under top-$k$ recommendation different $k$, on MovieLens 20M dataset
6 Conclusion
We define a dynamics model to capture the dual influence of recommender systems on user preferences and content creation. Despite being a simplification, we believe that our model captures the essence of a real-world recommender system, and our effort is an important initial endeavor to study diversity in recommender systems with dual influence. We theoretically and empirically demonstrate that, due to dual influence, myopically optimizing recommendation diversity could cause polarization and long-term diversity loss. And somewhat unexpectedly, designs not targeting diversity (e.g., top-$k$ recommendation) can alleviate polarization. The insights from our work are valuable to building healthy and sustainable recommender systems, and we believe our results can inspire more sophisticated solutions for improving the diversity of recommender systems to be developed.
Appendix A Additional Discussion on Related Works
Diversity in recommendations
Diversity, filter bubbles, and polarization in recommendations have been important research topics in recent years, and they are closely related concepts with different focuses. On the one hand, filter bubbles are frequently defined as decreasing recommendation diversity over time (Aridor et al., 2020), which describes both the process and the outcome of insufficiently diverse recommendations. On the other hand, polarization describes the negative outcome of insufficient mutual understanding between people (Santos et al., 2021). In content platforms, an example of polarization is people creating content with strong agreement or disagreement with other content under the same topic, e.g., political opinions. To combat these negative outcomes, previous works propose diversity-boosting approaches including re-ranking (Carbonell & Goldstein, 1998; Ziegler et al., 2005) and diversity-aware objective optimization (Su et al., 2013; Zhang & Hurley, 2008; Hurley, 2013; Wilhelm et al., 2018; Cheng et al., 2017; Zhang et al., 2023). Despite having positive effects in situations where user preferences and creation styles are fixed, these approaches overlooked the dynamic nature of recommender systems and our work shows that certain approaches will make long-term outcomes worse under the dual influence.
Opinion dynamics
Opinion dynamics study the effect of people exchanging opinions with others on social networks (Sarlette et al., 2007; Golub & Jackson, 2010; Li & Spong, 2014; Altafini & Lini, 2015). Our model of a recommender system with dual influence on users and creators resembles a bipartite social network, and our conclusion that the system converges to polarization is conceptually similar to people reaching consensus on social networks (Acemoğlu et al., 2013; Caponigro et al., 2015; Markdahl et al., 2018; Zhang et al., 2022). However, the technique we use to prove our conclusion (absorbing Markov chain) significantly differs from the main technique (stability of ODE) in the mentioned works.
Performative prediction
Works in performative prediction assume that predictive systems (e.g., recommender systems) impact the individuals interacting with those systems (e.g., users and creators) (Perdomo et al., 2020; Hardt et al., 2022). These impacts can be direct, such as individuals ostensibly modifying their features in order to obtain more desirable outcomes (Levanon & Rosenfeld, 2022). Prior works on the performative effects of recommender systems (Eilat & Rosenfeld, 2023; Dean & Morgenstern, 2022) only consider one-sided impact, either on users or on creators. Differing from them, our work focuses on two-sided impacts, i.e., on both users and creators.
Appendix B Experiments
We present experimental results on the behavior of the user-creator feature dynamics on synthetic data and real-world (MovieLens 20M) data and the effects of top-$k$ recommendation and truncation on the dynamics.
B.1 Synthetic Data Experiments
Setup
The dynamics is initialized by randomly generating user and creator features on the unit sphere in ${\mathbb{R}}^{d}$. We pick $d=10$, number of creators $n=50$, number of users $m=100$. We use the softmax recommendation probability function (2). We simulate the dynamics for $T=1000$ steps, repeated $100$ times each with a new initialization. We choose the sign impact function $g(\bm{u}_{j},\bm{v}_{i})=\mathrm{sign}(\langle\bm{u}_{j},\bm{v}_{i}\rangle)$ for creator updates. For user updates, we choose inner product $f(\bm{v}_{i},\bm{u}_{j})=\langle\bm{v}_{i},\bm{u}_{j}\rangle$. The inner product function is studied in previous works on users’ preference dynamics (but not creators’) (Dean & Morgenstern, 2022). Note that the inner product does not satisfy the condition $|f(\bm{v}_{i},\bm{u}_{j})|\geq L_{f}$ needed in Theorem 3.3. However, we still observe convergence to polarization in nearly all experiments. Thus, even when this condition does not hold, users and creators still tend towards polarization in practice.
Three key parameters in our model are $\beta$ (sensitivity of the softmax function), $\eta_{c}$ (creator update rate), and $\eta_{u}$ (user update rate). We set them to $\beta=1,\eta_{c}=\eta_{u}=0.1$, and change one parameter at a time to see its effect on the dynamics. We also test what happens when some dimensions of the user features are fixed features that are not updated.
Measures
To quantify the behavior of the dynamics, given user and creator feature vectors $(\bm{U},\bm{V})$ we compute the following measures, which cover diversity, relevancy, and polarization of the system:
  • Creator Diversity (CD): diversity of the creator features, measured by their average pairwise distance (Ziegler et al., 2005; Nguyen et al., 2014):
    $\mathrm{CD}(\bm{V})=\frac{1}{n(n-1)}\sum_{i=1}^{n}\sum_{j\neq i}\|\bm{v}_{i}-\bm{v}_{j}\|.$ (5)
  • Recommendation Diversity (RD): diversity of the contents recommended to a user, measured by the weighted variance of the contents:
    $\mathrm{RD}(\bm{U},\bm{V};\beta)=\frac{1}{m}\sum_{j=1}^{m}\sum_{i=1}^{n}p_{ij}\|\bm{v}_{i}-\overline{\bm{v}}_{j}\|^{2},$ (6)
    where $\overline{\bm{v}}_{j}=\sum_{i=1}^{n}p_{ij}\bm{v}_{i}$ and $p_{ij}=\frac{\exp(\beta\langle\bm{u}_{j},\bm{v}_{i}\rangle)}{\sum_{i=1}^{n}\exp(\beta\langle\bm{u}_{j},\bm{v}_{i}\rangle)}$.
  • Recommendation Relevance (RR): relevance of the contents recommended to a user, measured by the weighted average of inner products:
    $\mathrm{RR}(\bm{U},\bm{V};\beta)=\frac{1}{m}\sum_{j=1}^{m}\sum_{i=1}^{n}p_{ij}\langle\bm{u}_{j},\bm{v}_{i}\rangle.$ (7)
  • Tendency to Polarization (TP): This is a novel measure we propose to quantify how close the system is to consensus or bi-polarization, measured by the average absolute inner products between the creators:
    $\mathrm{TP}(\bm{V})=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{k=1}^{n}|\langle\bm{v}_{i},\bm{v}_{k}\rangle|.$ (8)
    $\mathrm{TP}(\bm{V})$ being closer to 1 means that the system is more polarized, because the term $|\langle\bm{v}_{i},\bm{v}_{k}\rangle|$ is $1$ iff the two vectors $\bm{v}_{i},\bm{v}_{k}$ are equal or opposite to each other.
It is worth noting that a high creator diversity is necessary for simultaneously achieving high recommendation relevance and high recommendation diversity. For example, they cannot be simultaneously achieved in a polarized state.
Sensitivity Parameter $\beta$
A larger $\beta$ means that a user will be recommended more relevant content/creator with a higher probability. $\beta=0$, on the other hand, means that the user receives uniform recommendations across all creators. Our main observation from the experiments is: a larger $\beta$ leads to higher creator diversity and alleviated polarization in the long run.
Figure 2 shows snapshots of the dynamics at different time steps under different $\beta$ values. Here, we choose dimension $d=3$ instead of $10$ so the feature vectors can be visualized on a 3d sphere. We see that the system tends to form more clusters at time $t=200$ as $\beta$ increases.
Figure 2: Snapshots of the dynamics simulated with the same initialization but different recommendation sensitivity $\beta$.
A larger $\beta$ resulted in more clusters at time step $t=200$.
Figure 3: Changes of measures over time under different sensitivity parameter $\beta$, on synthetic data
Figure 3 shows the changes of the 4 measures CD, RD, RR, TP over time under different $\beta$ values. We see that a more diverse recommendation policy (a smaller $\beta$) leads to lower creator diversity and a higher level of polarization in the long run. In particular, while Creator Diversity reaches a similar level under different $\beta$ in the end, it drops at a slower rate with a larger $\beta$ (see $\beta=5,6$). Moreover, from the plot of Tendency to Polarization, we see that a larger $\beta$ alleviates polarization, which means improvement in the diversity of the whole system.
An explanation for our observation is the following: when $\beta$ is smaller, each user receives more uniform recommendations across all creators, so for different creators the sets of users recommended to those creators have larger intersections. Since the creator updates are based on the sets of recommended users, different creators will be moving towards more similar directions. This leads to faster polarization. One can also predict this observation from Theorem 3.3: when $\beta$ is large, the minimum recommendation probability $p_{0}$ of the softmax function tends to 00, so it might take a long time for the system to converge to polarization, while with a small $\beta$ the system polarizes quickly.
Update Rates $\eta_{c}$ and $\eta_{u}$
A larger $\eta_{c}$ means that creator features are updated faster, and intuitively should lead to faster polarization. This is validated in experiments: Figure 4 shows that a larger $\eta_{c}$ indeed causes more extreme polarization and lower diversity (both CD and RD). A larger $\eta_{u}$ means that user features are updated faster. It has a similar effect of exacerbating polarization as $\eta_{c}$ does, as shown in Figure 5.
Figure 4: Changes of measures over time under different creator update rate $\eta_{c}$, on synthetic data
Figure 5: Changes of measures over time under different user update rate $\eta_{u}$, on synthetic data
Number of Fixed Dimensions
We also consider the scenario where some dimensions of the user feature vectors are fixed features and thus not updated from round to round (e.g., age, gender), which is a realistic scenario. Formally, we fix the first $k\leq d$ dimensions. The remaining $d-k$ dimensions $\bm{u}_{j}^{t}[k+1:d]=(u_{j}^{t}[k+1],\ldots,u_{j}^{t}[d])$ are updated according to the following rule: $\bm{u}_{j}^{t+1}[k+1:d]=\|\bm{u}_{j}^{t}[k+1:d]\|\cdot\mathcal{P}\big{(}\bm{u}_{j}^{t}[k+1:d]+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})\bm{v}_{i}^{t}[k+1:d]\big{)}$. The multiplication by $\|\bm{u}_{j}^{t}[k+1:d]\|$ ensures unit norm $\|\bm{u}_{j}^{t+1}\|=1$. The effect of the number of fixed dimensions $k$ on the dynamics is shown in Figure 6. The main observation is: as the number of fixed dimensions increases, the diversity of the system improves and the degree of polarization is reduced. This is similar to the effect of decreasing user update rate $\eta_{u}$. The observation that fixed dimensions of user features help to improve diversity might be a reason why the recommender systems in practice are not as polarized as our theoretical prediction.
Figure 6: Changes of measures over time under different numbers of fixed dimensions, on synthetic data
B.2 Real-World Data Experiments
Figure 7: Two tower model for the MovieLens experiment, where the two towers both have size $16\times 16$ with linear layers and ReLu activations.
Algorithm 1 Real-world Recommendation with Dual Influence
  Input: $t=0$, actual embedding $U^{(0)},V^{(0)}$, true labels $Y_{ij}^{(0)}:=y(u_{i}^{(0)},v_{j}^{(0)})$, initial parameter $\bm{\theta}^{(0)}$ (which includes the predicted embedding $\hat{U}^{(0)},\hat{V}^{(0)}$)
  repeat
     Let temporary parameter $\bm{w}^{(0)}\leftarrow\bm{\theta}^{(t)}$
     Compute loss $\mathcal{L}(\bm{\theta}^{(t)},Y^{(t)})$
     for $s=1$ to $m-1$ do
        $\bm{w}^{(s+1)}\leftarrow\bm{\theta}^{(s)}-\eta\nabla_{\bm{w}}\mathcal{L}(\bm{w}^{(s)},Y^{(t)})$
     end for
     $\bm{\theta}^{(t+1)}\leftarrow\bm{w}^{(m)}$
     Deliver recommendations based on $\hat{U}^{(t+1)},\hat{V}^{(t+1)}$
     Update $U^{(t+1)},V^{(t+1)}$, and $Y^{(t+1)}$
     $t\leftarrow t+1$
  until $\|\bm{\theta}^{(t)}-\bm{\theta}^{(t-1)}\|_{2}\leq\delta$
Figure 8: Experiment on MovieLens 20M dataset under different recommendation sensitivity $\beta$
Figure 9: Experiment on MovieLens 20M dataset with diversity-aware objective under different $\rho$
In this part, we conduct experiments on the MovieLens 20M dataset (Harper & Konstan, 2015). We use a real-world two-tower recommendation model with 16-dimensional tower tops as the user and creator embeddings (Fig. 7). The model was initialized on the original data, which we treat as the initial user and creator embeddings. Then we follow Algorithm 1 to simulate the dynamics.
Figure 8 shows the effect of the recommendation sensitivity parameter $\beta$ on the system. Similar to the synthetic data experiments, a smaller $\beta$ (more diverse recommendation for the users in the short term) results in faster polarization. We note that the joint results on CD and TP are more informative than each one alone: despite $\beta=0$ has a higher creator diversity than $\beta=2$ at $T=500$, the system reaches polarization more quickly under $\beta=0$. The higher creator diversity under $\beta=0$ is because the two clusters in the bi-polarized state are more balanced so the average pairwise distance between the creators is higher under $\beta=0$ than under $\beta=2$.
Figure 9 shows the effect of using diversity-aware objective for diversity boosting (see Section 4). We see that myopically promoting the short-term recommendation diversity (using a larger $\rho$) results in a higher tendency to polarization and a lower creation diversity in the long run.
B.3 Top-$k$ Recommendation
We experimented with top-$k$ recommendation on the synthetic data and the MovieLens 20M dataset. Our main observation is: a small $k$ improves the diversity of the recommender system and reduces polarization. This is consistent with our theoretical prediction (Proposition 4.1). However, there is a tradeoff between the diversity of recommendations to users (RD) and the diversity of creations in the system (CD and TP). A top-$k$ recommendation policy with small $k$ is “not diverse” for users because it exposes a user only to a small set of contents. However, such a policy can lead to a more diverse outcome in the whole system. This tradeoff is worth further studying.
Results on the synthetic data set are in Table 1.
Table 1: Diversity improvement by top-$k$ recommendation on synthetic data
$\beta$ $k$ Creator Diversity Recommendation Diversity Recommendation Relevance Tendency to Polarization
$1$ $50$ $1.00_{\pm.03}$ ${}~{}~{}\mathbf{0.42_{\pm 0.01}}$ ${}~{}0.76_{\pm 0.01}$ $1.00_{\pm 10^{-3}}$
$25$ $0.52_{\pm.32}$ ${}~{}0.03_{\pm 0.03}$ ${}~{}0.97_{\pm 0.02}$ $0.91_{\pm 0.13}$
$20$ $0.91_{\pm.15}$ ${}~{}0.00_{\pm 0.01}$ ${}~{}1.00_{\pm 0.01}$ $0.68_{\pm 0.12}$
$10$ $1.17_{\pm.06}$ $\quad 0.00_{\pm 10^{-3}}$ $\quad 1.00_{\pm 10^{-3}}$ $0.50_{\pm 0.07}$
$5$ $1.31_{\pm.02}$ $\quad 0.00_{\pm 10^{-3}}$ $\quad 1.00_{\pm 10^{-3}}$ $0.35_{\pm 0.03}$
$1$ $\quad\mathbf{1.40_{\pm 10^{-3}}}$ $\quad 0.00_{\pm 10^{-3}}$ ${}\quad\mathbf{1.00_{\pm 10^{-3}}}$ $\quad\mathbf{0.27_{\pm 10^{-3}}}$
$3$ $50$ $0.95_{\pm.14}$ ${}~{}~{}\mathbf{0.02_{\pm 0.02}}$ ${}~{}0.99_{\pm 0.01}$ $0.91_{\pm 0.10}$
$25$ $0.80_{\pm.24}$ ${}~{}0.00_{\pm 0.01}$ $\quad 1.00_{\pm 10^{-3}}$ $0.77_{\pm 0.13}$
$20$ $0.89_{\pm.13}$ $\quad 0.00_{\pm 10^{-3}}$ $\quad 1.00_{\pm 10^{-3}}$ $0.74_{\pm 0.11}$
$10$ $1.18_{\pm.05}$ $\quad 0.00_{\pm 10^{-3}}$ $\quad 1.00_{\pm 10^{-3}}$ $0.49_{\pm 0.07}$
$5$ $1.31_{\pm.02}$ $\quad 0.00_{\pm 10^{-3}}$ $\quad 1.00_{\pm 10^{-3}}$ $0.34_{\pm 0.03}$
$1$ $\quad\mathbf{1.40_{\pm 10^{-3}}}$ $\quad 0.00_{\pm 10^{-3}}$ ${}\quad\mathbf{1.00_{\pm 10^{-3}}}$ $\quad\mathbf{0.27_{\pm 10^{-3}}}$
For the experiments on MovieLens 20M dataset, we have $n=2000$ creators and $m=2000$ users, with feature dimension $d=16$. The results for top-$k$ recommendation are in Table 2 and Figure 10. Similar to the experiments with synthetic data, we see that a smaller $k$ improves Creator Diversity (CD) and Recommendation Relevance (RR), reduces Tendency to Polarization (TP), yet worsens Recommendation Diversity (RD).
Table 2: Diversity improvement by top-$k$ recommendation on MovieLens 20M dataset
$\beta$ $k$ CD RD RR TP
00 $2000$ $1.00\pm 10^{-3}$ $\mathbf{1.00\pm 10^{-3}}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$
$1000$ $0.30\pm 0.04$ $0.03\pm 0.01$ $0.88\pm 0.01$ $1.00\pm 10^{-3}$
$500$ $1.10\pm 0.06$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.43\pm 0.03$
$100$ $1.36\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.28\pm 0.01$
$10$ $1.40\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.20\pm 10^{-3}$
$1$ $\mathbf{1.40\pm 10^{-3}}$ $0.00\pm 10^{-3}$ $\mathbf{1.00\pm 10^{-3}}$ $\mathbf{0.20\pm 10^{-3}}$
$1$ $2000$ $1.00\pm 10^{-3}$ $\mathbf{0.42\pm 10^{-3}}$ $0.92\pm 0.01$ $1.00\pm 10^{-3}$
$1000$ $0.61\pm 0.16$ $0.03\pm 0.01$ $0.97\pm 0.01$ $0.90\pm 0.06$
$500$ $1.14\pm 0.04$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.41\pm 0.04$
$100$ $1.35\pm 0.01$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.27\pm 10^{-3}$
$10$ $1.40\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.20\pm 10^{-3}$
$1$ $\mathbf{1.40\pm 10^{-3}}$ $0.00\pm 10^{-3}$ $\mathbf{1.00\pm 10^{-3}}$ $\mathbf{0.20\pm 10^{-3}}$
$3$ $2000$ $0.92\pm 0.07$ $\mathbf{0.02\pm 0.01}$ $0.99\pm 10^{-3}$ $0.91\pm 0.05$
$1000$ $0.65\pm 0.18$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.69\pm 0.14$
$500$ $1.07\pm 0.07$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.48\pm 0.11$
$100$ $1.36\pm 0.01$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.27\pm 0.01$
$10$ $1.40\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.20\pm 10^{-3}$
$1$ $\mathbf{1.40\pm 10^{-3}}$ $0.00\pm 10^{-3}$ $\mathbf{1.00\pm 10^{-3}}$ $\mathbf{0.20\pm 10^{-3}}$
Figure 10: Changes of measures over time under different $k$, with $\beta=1$, on MovieLens 20M dataset
B.4 Truncation
We also experimented with truncation on synthetic data and MovieLens data. The effect of a large truncation threshold $\tau$ is similar to the effect of a small $k$ in top-$k$ recommendation.
Table 3 shows that the effect of truncation with different thresholds on synthetic data. We see that truncating at $\tau=0$, which corresponds to $90^{\circ}$ angle between $\bm{u}_{j}$ and $\bm{v}_{i}$, is not good for diversity, resulting in the lowest creator diversity measure (CD) and highest tendency to polarization (TP). Truncating at a large threshold like $0.707$ is good for diversity, instead. Figure 11 shows how the diversity measures change over time, under different truncation thresholds.
Results for truncation with different thresholds on the MovieLens 20M dataset are in Table 4 and Figure 12. Similar to synthetic data, we see that a large (but not too large) threshold like 0.707 is good for improving CD and TP.
Table 3: Inner product truncation with different thresholds on synthetic data
$\beta$ threshold $\tau$ CD RD RR TP
00 $-\cos(60^{\circ})=-0.5$ $1.00\pm 0.03$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.99\pm 10^{-3}$
$-\cos(72^{\circ})=-0.309$ $0.96\pm 0.06$ $\mathbf{0.01\pm 0.02}$ $1.00\pm 0.02$ $0.92\pm 0.10$
$\cos(90^{\circ})=0$ $0.03\pm 0.16$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.99\pm 0.04$
$\cos(72^{\circ})=0.309$ $0.72\pm 0.30$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.81\pm 0.12$
$\cos(60^{\circ})=0.5$ $1.16\pm 0.11$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.47\pm 0.10$
$\cos(45^{\circ})=0.707$ $\mathbf{1.37\pm 0.02}$ $0.00\pm 10^{-3}$ $\mathbf{1.00\pm 10^{-3}}$ $\mathbf{0.33\pm 0.02}$
$\cos(30^{\circ})=0.866$ $1.30\pm 0.03$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.55\pm 0.05$
$1$ $-\cos(60^{\circ})=-0.5$ $0.98\pm 0.04$ $0.00\pm 0.02$ $1.00\pm 0.01$ $0.96\pm 0.04$
$-\cos(72^{\circ})=-0.309$ $0.92\pm 0.08$ $0.00\pm 0.02$ $0.99\pm 0.02$ $0.87\pm 0.10$
$\cos(90^{\circ})=0$ $0.13\pm 0.31$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.97\pm 0.08$
$\cos(72^{\circ})=0.309$ $0.85\pm 0.16$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.76\pm 0.11$
$\cos(60^{\circ})=0.5$ $1.21\pm 0.07$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.43\pm 0.08$
$\cos(45^{\circ})=0.707$ $\mathbf{1.38\pm 0.01}$ $0.00\pm 10^{-3}$ $\mathbf{1.00\pm 10^{-3}}$ $\mathbf{0.30\pm 0.01}$
$\cos(30^{\circ})=0.866$ $1.33\pm 0.02$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.47\pm 0.04$
$3$ $-\cos(60^{\circ})=-0.5$ $0.91\pm 0.18$ $\mathbf{0.01\pm 0.02}$ $1.00\pm 0.01$ $0.83\pm 0.10$
$-\cos(72^{\circ})=-0.309$ $0.85\pm 0.23$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.78\pm 0.11$
$\cos(90^{\circ})=0$ $0.64\pm 0.33$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.81\pm 0.12$
$\cos(72^{\circ})=0.309$ $1.01\pm 0.14$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.64\pm 0.14$
$\cos(60^{\circ})=0.5$ $1.26\pm 0.05$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.38\pm 0.06$
$\cos(45^{\circ})=0.707$ $\mathbf{1.39\pm 0.01}$ $0.00\pm 10^{-3}$ $\mathbf{1.00\pm 10^{-3}}$ $\mathbf{0.28\pm 0.01}$
$\cos(30^{\circ})=0.866$ $1.37\pm 0.01$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.34\pm 0.01$
Figure 11: Changes of measures over time under different truncation threshold $\tau$, with $\beta=1$, on synthetic data
Table 4: Inner product truncation with different thresholds on MovieLens 20M dataset
$\beta$ threshold $\tau$ CD RD RR TP
00 $-\cos(60^{\circ})=-0.5$ $1.00\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $1.00\pm 10^{-3}$
$-\cos(72^{\circ})=-0.309$ $1.00\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $1.00\pm 10^{-3}$
$\cos(90^{\circ})=0$ $0.01\pm 0.01$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $1.00\pm 10^{-3}$
$\cos(72^{\circ})=0.309$ $0.83\pm 0.08$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.72\pm 0.09$
$\cos(60^{\circ})=0.5$ $1.20\pm 0.05$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.46\pm 0.07$
$\cos(45^{\circ})=0.707$ $\mathbf{1.39\pm 10^{-3}}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $\mathbf{0.20\pm 10^{-3}}$
$\cos(30^{\circ})=0.866$ $1.36\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.40\pm 0.01$
$1$ $-\cos(60^{\circ})=-0.5$ $1.00\pm 10^{-3}$ ${0.00\pm 10^{-3}}$ $1.00\pm 10^{-3}$ $1.00\pm 10^{-3}$
$-\cos(72^{\circ})=-0.309$ $0.96\pm 0.03$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.95\pm 0.03$
$\cos(90^{\circ})=0$ $0.02\pm 0.02$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.99\pm 10^{-3}$
$\cos(72^{\circ})=0.309$ $0.83\pm 0.07$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.66\pm 0.10$
$\cos(60^{\circ})=0.5$ $1.18\pm 0.06$ $0.00\pm 10^{-3}$ $1.00\pm 0.01$ $0.44\pm 0.07$
$\cos(45^{\circ})=0.707$ $\mathbf{1.40\pm 10^{-3}}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $\mathbf{0.20\pm 10^{-3}}$
$\cos(30^{\circ})=0.866$ $1.35\pm 0.01$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.40\pm 0.02$
$3$ $-\cos(60^{\circ})=-0.5$ $0.77\pm 0.27$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.86\pm 0.09$
$-\cos(72^{\circ})=-0.309$ $0.80\pm 0.24$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.79\pm 0.13$
$\cos(90^{\circ})=0$ $0.04\pm 0.02$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.98\pm 0.01$
$\cos(72^{\circ})=0.309$ $0.99\pm 0.11$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.55\pm 0.13$
$\cos(60^{\circ})=0.5$ $1.26\pm 0.05$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.36\pm 0.06$
$\cos(45^{\circ})=0.707$ $\mathbf{1.40\pm 10^{-3}}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $\mathbf{0.20\pm 10^{-3}}$
$\cos(30^{\circ})=0.866$ $1.36\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.39\pm 0.01$
Figure 12: Changes of measures over time under truncation with different threshold $\tau$, with $\beta=1$, on MovieLens 20M dataset
Appendix C Missing Proofs from Section 3
C.1 Proof of Proposition 3.2
We prove a stronger proposition than Proposition 3.2. We say that dynamics $(\bm{U}^{t},\bm{V}^{t})$ reach $(R,\bm{c})$-bi-polarization (or $(R,\bm{c})$-consensus) if the dynamics reaches $R$-bi-polarization (or $R$-consensus) with the vector $\bm{c}$. We prove that $(R,c)$-bi-polarization states are absorbing, which implies that $R$-bi-polarization states are absorbing.
Proof.
Let $(\bm{U}^{t},\bm{V}^{t})$ be an $(R,\bm{c})$-bi-polarization state with $R\in[0,1]$ and $\bm{c}\in\mathbb{S}^{d-1}$, where all $\bm{u}_{j}^{t}$ and $\bm{v}_{i}^{t}$ are within distance $R$ to $+\bm{c}$ or $-\bm{c}$. We show that, after one step of update, $\bm{u}_{j}^{t+1}$ and $\bm{v}_{i}^{t+1}$ are still within distance $R$ to $+\bm{c}$ or $-\bm{c}$, so $(\bm{U}^{t+1},\bm{V}^{t+1})$ still satisfies $(R,\bm{c})$-bi-polarization.
Consider $\bm{u}_{j}^{t}$. Without loss of generality, suppose $\bm{u}_{j}^{t}$ is close to $+\bm{c}$, so $\|\bm{u}_{j}^{t}-\bm{c}\|_{2}\leq R$. Suppose user $j$ is recommended creator $i$ at step $t$. Let $\tilde{\bm{v}}_{i}^{t}=\bm{v}_{i}^{t}$ if $\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}\rangle\geq 0$ and $\tilde{\bm{v}}_{i}^{t}=-\bm{v}_{i}^{t}$ if $\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}\rangle<0$. Then, the user update is
$\bm{u}_{j}^{t+1}=\mathcal{P}\Big{(}\bm{u}_{j}^{t}+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})\bm{v}_{i}^{t}\Big{)}=\mathcal{P}\Big{(}\bm{u}_{j}^{t}+\eta_{u}|f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})|\tilde{\bm{v}}_{i}^{t}\Big{)}.$
Since $\tilde{\bm{v}}_{i}^{t}$ is close to $+\bm{c}$ or $-\bm{c}$, $\langle\tilde{\bm{v}}_{i}^{t},\bm{u}_{j}^{t}\rangle>0$, and $\bm{u}_{j}^{t}$ is close to $+\bm{c}$, it must be that $\tilde{\bm{v}}_{i}^{t}$ is close to $+\bm{c}$, so $\|\tilde{\bm{v}}_{i}^{t}-\bm{c}\|_{2}\leq R$. Then, since $\bm{u}_{j}^{t+1}$ is the normalization of a vector in the convex cone formed by $\bm{u}_{j}^{t}$ and $\tilde{\bm{v}}_{i}^{t}$, by Lemma E.2, we have
$\|\bm{u}_{j}^{t+1}-\bm{c}\|_{2}~{}\leq~{}\max\big{\{}\|\bm{u}_{j}^{t}-\bm{c}\|_{2},~{}\|\tilde{\bm{v}}_{i}^{t}-\bm{c}\|_{2}\big{\}}~{}\leq~{}R.$
Consider $\bm{v}_{i}^{t}$. Suppose $\|\bm{v}_{i}^{t}-\bm{c}\|_{2}\leq R$. Let $J$ be the set of users that are recommended creator $i$ at step $t$. For each $j\in J$, let $\tilde{\bm{u}}_{j}^{t}=\bm{u}_{j}^{t}$ if $\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\geq 0$ and $\tilde{\bm{u}}_{j}^{t}=-\bm{u}_{j}^{t}$ if $\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle<0$. Then, the creator update is
$\bm{v}_{i}^{t+1}$ $=\mathcal{P}\Big{(}\bm{v}_{i}^{t}+\frac{\eta_{c}}{|J|}\sum_{j\in J}g(\bm{u}_{j}^{t},\bm{v}_{i}^{t})\bm{u}_{j}^{t}\Big{)}=\mathcal{P}\Big{(}\bm{v}_{i}^{t}+\frac{\eta_{c}}{|J|}\sum_{j\in J}|g(\bm{u}_{j}^{t},\bm{v}_{i}^{t})|\tilde{\bm{u}}_{j}^{t}\Big{)}.$
We note that every $\tilde{\bm{u}}_{j}^{t}$ satisfies $\|\tilde{\bm{u}}_{j}^{t}-\bm{c}\|_{2}\leq R$ (by the same reasoning as above). Then, since $\bm{v}_{i}^{t+1}$ is the normalization of a vector in the convex cone formed by $\bm{v}_{i}^{t}$ and $\{\tilde{\bm{u}}_{j}^{t}\}_{j\in J}$, by Lemma E.2, we have
$\|\bm{v}_{i}^{t+1}-\bm{c}\|_{2}~{}\leq~{}\min\Big{\{}\|\bm{v}_{i}^{t}-\bm{c}\|_{2},~{}\min_{j\in J}\|\tilde{\bm{u}}_{j}^{t}-\bm{c}\|_{2}\Big{\}}~{}\leq~{}R.\qed$
C.2 Proof of Theorem 3.3
The proof of Theorem 3.3 uses the following lemma:
Lemma C.1.
Suppose $\eta_{c}\leq\frac{\eta_{u}L_{f}}{2}$ and $\eta_{u}<\frac{1}{2}$. For any $R>0$, for almost every state $(\bm{U}^{t},\bm{V}^{t})$ in the state space, there exists a path $(\bm{U}^{t},\bm{V}^{t})\to(\bm{U}^{t+1},\bm{V}^{t+1})\to\cdots\to(\bm{U}^{t+T},\bm{V}^{t+T})$ of finite length that leads to an $R$-bi-polarization state $(\bm{U}^{t+T},\bm{V}^{t+T})$.
The proof of this lemma (in Appendix F) is involved. It uses induction on the number of creators $n$. The base case of $n=1$ is proved by a potential function argument. For $n\geq 2$, we first construct a path that leads the subsystem of $n-1$ creators and all users to $R$-bi-polarization. Then, depending on where the remaining creator is, we construct a sequence of recommendations that leads the remaining creator to one of the two clusters formed by the $n-1$ creators and all users. Such recommendations will move some users out of the formed clusters, which requires extra care in the proof.
Proof of Theorem 3.3.
For any state $(\bm{U}^{t},\bm{V}^{t})$ in the state space, by Lemma C.1 there exists a path $(\bm{U}^{t},\bm{V}^{t})\to\cdots\to(\bm{U}^{t+T},\bm{V}^{t+T})$ of length $T$ that leads to $R$-bi-polarization. Because every creator can be recommended to a user with probability at least $p_{0}$, each transition $(\bm{U}^{t^{\prime}},\bm{V}^{t^{\prime}})\to(\bm{U}^{t^{\prime}+1},\bm{V}^{t^{\prime}+1})$ happens with probability at least $p_{0}^{m}$. So, the path of length $T$ has probability at least $p_{0}^{mT}>0$, and the probability that the dynamics does not reach $R$-bi-polarization after $KT$ steps is at most $(1-p_{0}^{mT})^{K}$, which $\to 0$ as $K\to\infty$. Therefore, with probability $1$ the dynamics will reach $R$-bi-polarization eventually. ∎
Appendix D Missing Proofs from Section 4
D.1 Proof of Proposition 4.1
Before proving Proposition 4.1, we formally define what “clusters” mean:
Definition D.1 (clusters).
We say a state $(\bm{U}^{t},\bm{V}^{t})$ forms $q$ clusters if there exist $\bm{c}_{1},\ldots,\bm{c}_{q}\in{\mathbb{R}}^{d}$ and a small number $R>0$ such that every feature vector is in the $\ell_{2}$ ball of some $\bm{c}_{i}$ with radius $R$ (denoted by $B(\bm{c}_{\ell},R)=\{\bm{x}:\|\bm{x}-\bm{c}_{\ell}\|_{2}\leq R\}$), and $B(\bm{c}_{\ell},2R)\cap B(\bm{c}_{\ell^{\prime}},2R)=\emptyset$ for $\ell\neq\ell^{\prime}$.
It is clear that consensus has a single cluster, and bi-polarization has two.
Proof of Proposition 4.1.
Let $R>0$ be any small number. Let $\bm{c}_{1},\ldots,\bm{c}_{\lfloor n/k\rfloor}\in{\mathbb{R}}^{d}$ be $\lfloor n/k\rfloor$ vectors that satisfy $B(\bm{c}_{\ell},2R)\cap B(\bm{c}_{\ell^{\prime}},2R)=\emptyset$ for $\ell\neq\ell^{\prime}$, where $B(\bm{c},R)$ is the ball centered at $\bm{c}$ with radius $R$: $\{\bm{x}\in{\mathbb{R}}^{d}:\|\bm{x}-\bm{c}\|_{2}\leq R\}$. Consider user and creator features $(\bm{U}^{t},\bm{V}^{t})$ that satisfy: every ball $B(\bm{c}_{\ell},R)$ ($\ell=1,\ldots,\lfloor n/k\rfloor$) contains $k$ creator vectors, and every user vector $\bm{u}_{j}^{t}$ is in one of the balls $B(\bm{c}_{\ell},R)$. By definition, $(\bm{U}^{t},\bm{V}^{t})$ form $\lfloor n/k\rfloor$ clusters. We show that, after one step of update, the new state $(\bm{U}^{t+1},\bm{V}^{t+1})$ must still form $\lfloor n/k\rfloor$ clusters. Consider any user $j$. Suppose $\bm{u}_{j}^{t}\in B(\bm{c}_{\ell},R)$, then the distance from $\bm{u}_{j}^{t}$ to any creator $\bm{v}_{i}^{t}\in B(\bm{c}_{\ell},R)$ is at most $2R$:
$\|\bm{u}_{j}^{t}-\bm{v}_{i}^{t}\|\leq 2R.$
The distance from $\bm{u}_{j}^{t}$ to any creator $\bm{v}_{i^{\prime}}^{t}$ not in $B(\bm{c}_{\ell},R)$ is greater than $2R$:
$\|\bm{u}_{j}^{t}-\bm{v}_{i^{\prime}}^{t}\|>2R$
because $\bm{v}_{i^{\prime}}^{t}$ is in some other ball $B(\bm{c}_{\ell^{\prime}},R)$ that satisfies $B(\bm{c}_{\ell^{\prime}},2R)\cap B(\bm{c}_{\ell},2R)=\emptyset$. This implies that the inner products between user $j$ and the creators in ball $B(\bm{c}_{\ell},R)$ are greater than that with the creators in other ball:
$\forall\bm{v}_{i}^{t}\in B(\bm{c}_{\ell},R),~{}~{}\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle=1-\frac{1}{2}\|\bm{u}_{j}^{t}-\bm{v}_{i}^{t}\|_{2}^{2}$ $\geq 1-\frac{1}{2}(2R)^{2}>1-\frac{1}{2}\|\bm{u}_{j}^{t}-\bm{v}_{i^{\prime}}^{t}\|=\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle,~{}~{}\forall\bm{v}_{i^{\prime}}^{t}\in B(\bm{c}_{\ell^{\prime}},R).$
Since $B(\bm{c}_{\ell},R)$ contains $k$ creators, these $k$ creators are the $k$-most relevant ones to user $j$, so user $j$ will only be recommended these creators. Then, by applying Proposition 3.2 to each of the $\lfloor n/k\rfloor$ balls separately, we see that each ball is a $R$-consensus and hence absorbing. So, the new state $(\bm{U}^{t+1},\bm{V}^{t+1})$ still forms $\lfloor n/k\rfloor$ clusters with these $\lfloor n/k\rfloor$ balls. ∎
Appendix E Useful Lemmas
This section provides some lemmas that will be used in the proofs. They are mainly about some properties of the dynamics update rule.
Claim E.1.
For vectors $\bm{x},\bm{y}\in{\mathbb{R}}^{d}$ with unit norm $\|\bm{x}\|_{2}=\|\bm{y}\|_{2}=1$, we have:
  • $\|\bm{x}-\bm{y}\|_{2}^{2}=2(1-\langle\bm{x},\bm{y}\rangle)$.
  • $\langle\bm{x},\bm{y}\rangle=1-\frac{1}{2}\|\bm{x}-\bm{y}\|_{2}^{2}$.
Lemma E.2 (Convex Cone Property).
Let $\bm{z}_{1},\ldots,\bm{z}_{k}\in{\mathbb{R}}^{d}$ be vectors with norm $\|\bm{z}_{i}^{t}\|_{2}=1$. Suppose $\langle\bm{z}_{i},\bm{y}\rangle>0$ for every $i=1,\ldots,k$ for some $\bm{y}\in{\mathbb{R}}^{d}$. Let $\bm{x}=\mathcal{P}(\sum_{i=1}^{k}a_{i}\bm{z}_{i})$ for some $a_{1},\ldots,a_{k}\geq 0$ (namely, $\bm{x}$ is the normalization of some vector in the convex cone formed by $\bm{z}_{1},\ldots,\bm{z}_{k}$). Then, we have
$\langle\bm{x},\bm{y}\rangle~{}\geq~{}\min_{i=1}^{k}\langle\bm{z}_{i},\bm{y}\rangle~{}>~{}0\quad\text{ and }\quad\|\bm{x}-\bm{y}\|_{2}~{}\leq~{}\max_{i=1}^{k}\|\bm{z}_{i}-\bm{y}\|_{2}~{}>~{}0.$
Proof.
$\langle\bm{x},\,\bm{y}\rangle$ $~{}=~{}\Big{\langle}\frac{\sum_{i=1}^{k}a_{i}\bm{z}_{i}}{\|\sum_{i=1}^{k}a_{i}\bm{z}_{i}\|_{2}},\,\bm{y}\Big{\rangle}~{}=~{}\frac{1}{\|\sum_{i=1}^{k}a_{i}\bm{z}_{i}\|_{2}}\sum_{i=1}^{k}a_{i}\langle\bm{z}_{i},\bm{y}\rangle$
$~{}\geq~{}\frac{1}{\|\sum_{i=1}^{k}a_{i}\bm{z}_{i}\|_{2}}\sum_{i=1}^{k}a_{i}\min_{i=1}^{k}\langle\bm{z}_{i},\bm{y}\rangle~{}=~{}\min_{i=1}^{k}\langle\bm{z}_{i},\bm{y}\rangle\frac{\sum_{i=1}^{k}a_{i}}{\|\sum_{i=1}^{k}a_{i}\bm{z}_{i}\|_{2}}$
$~{}\geq~{}\min_{i=1}^{k}\langle\bm{z}_{i},\bm{y}\rangle\frac{\sum_{i=1}^{k}a_{i}}{\sum_{i=1}^{k}a_{i}}~{}=~{}\min_{i=1}^{k}\langle\bm{z}_{i},\bm{y}\rangle.$
This proves the first inequality. To prove the second inequality, we use Claim E.1 and the first inequality:
$\|\bm{x}-\bm{y}\|_{2}~{}=~{}\sqrt{2(1-\langle\bm{x},\bm{y}\rangle)}~{}\leq~{}\sqrt{2(1-\min_{i}\langle\bm{z}_{i},\bm{y}\rangle)}~{}=~{}\max_{i}\sqrt{2(1-\langle\bm{z}_{i},\bm{y}\rangle)}~{}=~{}\max_{i=1}^{k}\|\bm{z}_{i}-\bm{y}\|_{2}.$
Lemma E.3.
Let $\bm{x}^{t},\bm{y},\bm{z}^{t}\in{\mathbb{R}}^{d}$ be vectors with norm $\|\bm{x}^{t}\|_{2}=1$, $\|\bm{y}\|_{2}\geq 0$, $\|\bm{z}^{t}\|_{2}\leq 1$. Suppose $\langle\bm{x}^{t},\bm{y}\rangle\geq 0$, $\langle\bm{z}^{t},\bm{y}\rangle\geq 0$. After the update $\bm{x}^{t+1}=\mathcal{P}(\bm{x}^{t}+\eta\bm{z}^{t})$, we have
$\langle\bm{x}^{t+1}-\bm{x}^{t},\,\bm{y}\rangle~{}\geq~{}\frac{\eta}{1+\eta\|\bm{z}^{t}\|_{2}}\Big{(}\langle\bm{z}^{t},\bm{y}\rangle-\|\bm{z}^{t}\|_{2}\langle\bm{x}^{t},\bm{y}\rangle\Big{)}.$
As a corollary, if $\bm{y}=\bm{z}^{t}$ and $\|\bm{z}^{t}\|_{2}=1$, then
$\langle\bm{x}^{t+1}-\bm{x}^{t},\,\bm{z}^{t}\rangle~{}\geq~{}\frac{\eta}{1+\eta}\Big{(}1-\langle\bm{x}^{t},\bm{z}^{t}\rangle\Big{)}.$
Proof.
By definition,
$\langle\bm{x}^{t+1}-\bm{x}^{t},\,\bm{y}\rangle$ $~{}=~{}\Big{\langle}\frac{\bm{x}^{t}+\eta\bm{z}^{t}}{\|\bm{x}^{t}+\eta\bm{z}^{t}\|_{2}}-\bm{x}^{t},\,\bm{y}\Big{\rangle}$
$~{}=~{}\Big{(}\frac{1}{\|\bm{x}^{t}+\eta\bm{z}^{t}\|_{2}}-1\Big{)}\cdot\langle\bm{x}^{t},\bm{y}\rangle\,+\,\frac{\eta}{\|\bm{x}^{t}+\eta\bm{z}^{t}\|_{2}}\cdot\langle\bm{z}^{t},\bm{y}\rangle$
(because ${\small(\text{because $\|\bm{x}^{t}+\eta\bm{z}^{t}\|_{2}\leq 1+\eta\|\bm{z}^{t}\|_{2}$})}$ $~{}\geq~{}\Big{(}\frac{1}{1+\eta\|\bm{z}^{t}\|_{2}}-1\Big{)}\cdot\langle\bm{x}^{t},\bm{y}\rangle\,+\,\frac{\eta}{1+\eta\|\bm{z}^{t}\|_{2}}\cdot\langle\bm{z}^{t},\bm{y}\rangle$
$~{}=~{}\frac{\eta}{1+\eta\|\bm{z}^{t}\|_{2}}\Big{(}\langle\bm{z}^{t},\bm{y}\rangle-\|\bm{z}^{t}\|_{2}\langle\bm{x}^{t},\bm{y}\rangle\Big{)}.$
Lemma E.4.
Let $\bm{x}^{t},\bm{z}^{t}\in{\mathbb{R}}^{d}$ be vectors with norm $\|\bm{x}^{t}\|_{2}=1$, $\|\bm{z}^{t}\|_{2}\leq 1$. Suppose $\langle\bm{x}^{t},\bm{z}^{t}\rangle\geq 0$ and $\eta>0$. Then the update $\bm{x}^{t+1}=\mathcal{P}(\bm{x}^{t}+\eta\bm{z}^{t})$ satisfies
  • $\langle\bm{x}^{t+1}-\bm{x}^{t},\bm{z}^{t}\rangle\,\geq\,\frac{1}{\eta}\|\bm{x}^{t+1}-\bm{x}^{t}\|_{2}^{2}$.
  • $\|\bm{x}^{t+1}-\bm{x}^{t}\|_{2}\,\leq\,\eta\|\bm{z}^{t}\|_{2}$.
Proof.
Let $\tilde{\bm{x}}^{t+1}=\bm{x}^{t}+\eta\bm{z}^{t}$, so $\bm{x}^{t}=\mathcal{P}(\tilde{\bm{x}}^{t+1})$ and $\bm{z}^{t}=\frac{1}{\eta}(\tilde{\bm{x}}^{t+1}-\bm{x}^{t})$. Then we have
$\langle\bm{x}^{t+1}-\bm{x}^{t},\,\bm{z}^{t}\rangle~{}=~{}\frac{1}{\eta}\langle\bm{x}^{t+1}-\bm{x}^{t},\,\tilde{\bm{x}}^{t+1}-\bm{x}^{t}\rangle.$
Because $\langle\bm{x}^{t},\bm{z}^{t}\rangle\geq 0$, the vector $\tilde{\bm{x}}^{t+1}=\bm{x}^{t}+\eta\bm{z}^{t}$ has length $\geq 1$ and hence is outside (or on the surface) of the $d$-dimensional unit ball. Since $\bm{x}^{t}=\mathcal{P}(\tilde{\bm{x}}^{t+1})$ is the projection of $\tilde{\bm{x}}^{t+1}$ onto the unit ball, and $\bm{z}^{t}$ is another vector inside the unit ball, by the “Pythagorean property” (Proposition 2.2 in (Bansal & Gupta, 2019)), we must have $\langle\bm{x}^{t}-\bm{x}^{t+1},\tilde{\bm{x}}^{t+1}-\bm{x}^{t+1}\rangle\leq 0$. This implies
$\langle\bm{x}^{t+1}-\bm{x}^{t},\,\bm{z}^{t}\rangle$ $~{}\geq~{}\frac{1}{\eta}\Big{(}\langle\bm{x}^{t+1}-\bm{x}^{t},\,\tilde{\bm{x}}^{t+1}-\bm{x}^{t}\rangle+\langle\bm{x}^{t}-\bm{x}^{t+1},\,\tilde{\bm{x}}^{t+1}-\bm{x}^{t+1}\rangle\Big{)}$
$~{}=~{}\frac{1}{\eta}\langle\bm{x}^{t+1}-\bm{x}^{t},\,\bm{x}^{t+1}-\bm{x}^{t}\rangle~{}=~{}\frac{1}{\eta}\|\bm{x}^{t+1}-\bm{x}^{t}\|_{2}^{2},$
which proves the first claim. To prove the second claim, we use Cauchy-Schwarz inequality:
$\frac{1}{\eta}\|\bm{x}^{t+1}-\bm{x}^{t}\|_{2}^{2}~{}\leq~{}\langle\bm{x}^{t+1}-\bm{x}^{t},\,\bm{z}^{t}\rangle~{}\leq~{}\|\bm{x}^{t+1}-\bm{x}^{t}\|_{2}\|\bm{z}^{t}\|_{2}.$
This implies $\|\bm{x}^{t+1}-\bm{x}^{t}\|_{2}\leq\eta\|\bm{z}^{t}\|_{2}$. ∎
Lemma E.5.
Consider a creator $\bm{v}_{i}^{t}$ and a user $\bm{u}_{j}^{t}$. Suppose the user is always recommended creator $i$ (so the user is updated by $\bm{u}_{j}^{t+1}=\mathcal{P}(\bm{u}_{j}^{t}+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})\bm{v}_{i}^{t})$), and creator $i$ is updated by $\bm{v}_{i}^{t+1}=\mathcal{P}(\bm{v}_{i}^{t}+\eta_{c}\bm{\alpha}_{i}^{t})$ with $\|\bm{\alpha}_{i}^{t}\|_{2}\leq 1$ and $\langle\bm{v}_{i}^{t},\bm{\alpha}_{i}^{t}\rangle\geq 0$ at each time step. Assume:
  • The inner product $\langle\bm{u}_{j}^{0},\bm{v}_{i}^{0}\rangle>0$ initially. (Note that $\langle\bm{u}_{j}^{0},\bm{u}_{j^{\prime}}^{0}\rangle$ needs not hold.)
  • There exists some constant $L_{f}>0$ such that $f(\bm{v}_{i},\bm{u}_{j})\geq L_{f}>0$ whenever $\langle\bm{u}_{j},\bm{v}_{i}\rangle>0$.
  • $\eta_{c}\leq\frac{\eta_{u}L_{f}}{2}$ and $0\leq\eta_{u}<\frac{1}{2}$.
Then, we have $\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle>0$ in all time steps.
Proof.
We prove by induction. Suppose $\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle>0$ already holds. We prove that $\langle\bm{u}_{j}^{t+1},\bm{v}_{i}^{t+1}\rangle>0$ will also hold. Take the difference between $\langle\bm{u}_{j}^{t+1},\bm{v}_{i}^{t+1}\rangle$ and $\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle$:
$\langle\bm{u}_{j}^{t+1},\bm{v}_{i}^{t+1}\rangle-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\,=\,\langle\bm{u}_{j}^{t+1},\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t}\rangle+\langle\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle.$
For $\langle\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle$, using Lemma E.3 with $\bm{x}^{t}=\bm{u}_{j}^{t},\bm{z}^{t}=\bm{v}_{i}^{t}$, and $\eta=\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})$, we get
$\langle\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle~{}\geq~{}\tfrac{\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})}{1+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}~{}\geq~{}\tfrac{\eta_{u}L_{f}}{1+\eta_{u}L_{f}}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}.$
For $\langle\bm{u}_{j}^{t+1},\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t}\rangle$, by Cauchy-Schwarz inequality and Lemma E.4,
$\langle\bm{u}_{j}^{t+1},\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t}\rangle$ $~{}\geq~{}-\|\bm{u}_{j}^{t+1}\|_{2}\cdot\|\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t}\|_{2}~{}\geq~{}-1\cdot\eta_{c}\|\bm{\alpha}_{i}^{t}\|_{2}~{}\geq~{}-\eta_{c}.$
  • If $1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle>\frac{1}{2}(1+\eta_{u}L_{f})$, then we have
    $\langle\bm{u}_{j}^{t+1},\bm{v}_{i}^{t+1}\rangle-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle~{}>~{}\eta_{u}L_{f}\tfrac{1}{2}-\eta_{c}~{}\geq 0$
    by the assumption of $\eta_{c}\leq\frac{\eta_{u}L_{f}}{2}$.
  • If $1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\leq\frac{1}{2}(1+\eta_{u}L_{f})$, then we have
    $\langle\bm{u}_{j}^{t+1},\bm{v}_{i}^{t+1}\rangle-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle~{}\geq~{}0-\eta_{c}$
[…]
    under the assumption of $\eta_{c}\leq\frac{\eta_{u}L_{f}}{2}$ and $\eta_{u}<\frac{1}{2}$.
The above two cases together ensure $\langle\bm{u}_{j}^{t+1},\bm{v}_{i}^{t+1}\rangle>0$. ∎
Lemma E.6.
Consider a system of one user and one creator that satisfies $\langle\bm{u}_{j}^{0},\bm{v}_{i}^{0}\rangle>0$ and $\langle\bm{u}_{j}^{0},\bm{y}\rangle>\langle\bm{v}_{i}^{0},\bm{y}\rangle>0$ for some $\bm{y}\in{\mathbb{R}}^{d}$ with $\|\bm{y}\|\leq 1$ initially. The creator is always recommended to the user (so the updates are $\bm{u}_{j}^{t+1}=\mathcal{P}(\bm{u}_{j}^{t}+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})\bm{v}_{i}^{t})$ and $\bm{v}_{i}^{t+1}=\mathcal{P}(\bm{v}_{i}^{t}+\eta_{c}\bm{u}_{j}^{t})$). Suppose $\eta_{c}\leq\frac{\eta_{u}L_{f}}{2}$ and $0\leq\eta_{u}<\frac{1}{2}$. Then, we have:
  • $\langle\bm{u}_{j}^{t},\bm{y}\rangle>\langle\bm{v}_{i}^{t},\bm{y}\rangle>0$ for all $t\geq 1$.
  • Suppose $\langle\bm{u}_{j}^{0},\bm{y}\rangle-\langle\bm{v}_{i}^{0},\bm{y}\rangle=D>0$. For any $R<D$, after $T=\frac{8}{3\eta_{u}L_{f}}\ln\frac{2}{R^{2}}$ steps, we have $\langle\bm{v}_{i}^{T},\bm{y}\rangle-\langle\bm{v}_{i}^{0},\bm{y}\rangle\geq\frac{\eta_{c}}{\eta_{u}+\eta_{c}}(D-R)$.
Proof.
We prove the first item by induction. Suppose $\langle\bm{u}_{j}^{t},\bm{y}\rangle>\langle\bm{v}_{i}^{t},\bm{y}\rangle>0$ holds. Consider $t+1$. First, by Lemma E.2, $\langle\bm{v}_{i}^{t+1},\bm{y}\rangle>0$ holds. Then, we prove $\langle\bm{u}_{j}^{t+1},\bm{y}\rangle>\langle\bm{v}_{i}^{t+1},\bm{y}\rangle$. Let $f=f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})$.
$\langle\bm{u}_{j}^{t+1},\bm{y}\rangle-\langle\bm{v}_{i}^{t+1},\bm{y}\rangle~{}=~{}\Big{\langle}\frac{\bm{u}_{j}^{t}+\eta_{u}f\bm{v}_{i}^{t}}{\|\bm{u}_{j}^{t}+\eta_{u}f\bm{v}_{i}^{t}\|_{2}},\bm{y}\Big{\rangle}-\Big{\langle}\frac{\bm{v}_{i}^{t}+\eta_{c}\bm{u}_{j}^{t}}{\|\bm{v}_{i}^{t}+\eta_{c}\bm{u}_{j}^{t}\|_{2}},\bm{y}\Big{\rangle}$
$~{}=~{}\Big{(}\frac{1}{\|\bm{u}_{j}^{t}+\eta_{u}f\bm{v}_{i}^{t}\|_{2}}-\frac{\eta_{c}}{\|\bm{v}_{i}^{t}+\eta_{c}\bm{u}_{j}^{t}\|_{2}}\Big{)}\langle\bm{u}_{j}^{t},\bm{y}\rangle-\Big{(}\frac{1}{\|\bm{v}_{i}^{t}+\eta_{c}\bm{u}_{j}^{t}\|_{2}}-\frac{\eta_{u}f}{\|\bm{u}_{j}^{t}+\eta_{u}f\bm{v}_{i}^{t}\|_{2}}\Big{)}\langle\bm{v}_{i}^{t},\bm{y}\rangle$
$~{}>~{}\Big{(}\frac{1}{\|\bm{u}_{j}^{t}+\eta_{u}f\bm{v}_{i}^{t}\|_{2}}-\frac{\eta_{c}}{\|\bm{v}_{i}^{t}+\eta_{c}\bm{u}_{j}^{t}\|_{2}}\Big{)}\langle\bm{v}_{i}^{t},\bm{y}\rangle-\Big{(}\frac{1}{\|\bm{v}_{i}^{t}+\eta_{c}\bm{u}_{j}^{t}\|_{2}}-\frac{\eta_{u}f}{\|\bm{u}_{j}^{t}+\eta_{u}f\bm{v}_{i}^{t}\|_{2}}\Big{)}\langle\bm{v}_{i}^{t},\bm{y}\rangle$
$~{}=~{}\Big{(}\frac{1+\eta_{u}f}{\|\bm{u}_{j}^{t}+\eta_{u}f\bm{v}_{i}^{t}\|_{2}}-\frac{1+\eta_{c}}{\|\bm{v}_{i}^{t}+\eta_{c}\bm{u}_{j}^{t}\|_{2}}\Big{)}\langle\bm{v}_{i}^{t},\bm{y}\rangle$
$~{}=~{}\Big{(}\frac{1+\eta_{u}f}{\sqrt{1+2\eta_{u}f\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle+(\eta_{u}f)^{2}}}-\frac{1+\eta_{c}}{\sqrt{1+2\eta_{c}\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle+(\eta_{c})^{2}}}\Big{)}\langle\bm{v}_{i}^{t},\bm{y}\rangle.$
Let $a=\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\leq 1$. We note that the function
$h(\eta)=\frac{1+\eta}{\sqrt{1+2\eta a+\eta^{2}}}=\sqrt{\frac{1+2\eta+\eta^{2}}{1+2\eta a+\eta^{2}}}=\sqrt{1+\frac{(2-2a)\eta}{1+2\eta a+\eta^{2}}}=\sqrt{1+\frac{2(1-a)}{\frac{1}{\eta}+2a+\eta}}$
is increasing in $\eta\in[0,1]$. Under the assumption of $\eta_{c}\leq\frac{\eta_{u}L_{f}}{2}\leq\frac{\eta_{u}f}{2}<\eta_{u}f$, we have $h(\eta_{c})\leq h(\eta_{u}f)$ and hence
$\langle\bm{u}_{j}^{t+1},\bm{y}\rangle-\langle\bm{v}_{i}^{t+1},\bm{y}\rangle~{}>~{}\big{(}h(\eta_{u}f)-h(\eta_{c})\big{)}\langle\bm{v}_{i}^{t},\bm{y}\rangle~{}\geq~{}0.$
We then prove the second item. Using Lemma E.3 for $\bm{v}_{i}^{t+1}=\mathcal{P}(\bm{v}_{i}^{t}+\eta_{c}\bm{u}_{j}^{t})$, we get
$\langle\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t},\bm{y}\rangle\geq\frac{\eta_{c}}{1+\eta_{c}}\Big{(}\langle\bm{u}_{j}^{t},\bm{y}\rangle-\langle\bm{v}_{i}^{t},\bm{y}\rangle\Big{)}.$
Using Lemma E.3 for $\bm{u}_{j}^{t+1}=\mathcal{P}(\bm{u}_{j}^{t}+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})\bm{v}_{i}^{t})$ and using the fact $\langle\bm{v}_{i}^{t},\bm{y}\rangle-\langle\bm{u}_{j}^{t},\bm{y}\rangle<0$ proved in item 1,
$\langle\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t},\bm{y}\rangle\geq\frac{\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})}{1+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})}\Big{(}\langle\bm{v}_{i}^{t},\bm{y}\rangle-\langle\bm{u}_{j}^{t},\bm{y}\rangle\Big{)}\geq\frac{\eta_{u}}{1+\eta_{u}}\Big{(}\langle\bm{v}_{i}^{t},\bm{y}\rangle-\langle\bm{u}_{j}^{t},\bm{y}\rangle\Big{)}.$
Rearranging the above two inequalities:
$\frac{1+\eta_{c}}{\eta_{c}}\Big{(}\langle\bm{v}_{i}^{t+1},\bm{y}\rangle-\langle\bm{v}_{i}^{t},\bm{y}\rangle\Big{)}\geq\langle\bm{u}_{j}^{t},\bm{y}\rangle-\langle\bm{v}_{i}^{t},\bm{y}\rangle;$
$\frac{1+\eta_{u}}{\eta_{u}}\Big{(}\langle\bm{u}_{j}^{t+1},\bm{y}\rangle-\langle\bm{u}_{j}^{t},\bm{y}\rangle\Big{)}\geq\langle\bm{v}_{i}^{t},\bm{y}\rangle-\langle\bm{u}_{j}^{t},\bm{y}\rangle.$
Summing the above two inequalities over $t=0,1,\ldots,T-1$:
$\frac{1+\eta_{c}}{\eta_{c}}\Big{(}\langle\bm{v}_{i}^{T},\bm{y}\rangle-\langle\bm{v}_{i}^{0},\bm{y}\rangle\Big{)}+\frac{1+\eta_{u}}{\eta_{u}}\Big{(}\langle\bm{u}_{j}^{T},\bm{y}\rangle-\langle\bm{u}_{j}^{0},\bm{y}\rangle\Big{)}\geq 0.$ (9)
According to Lemma F.1, after at most $T=\frac{8}{3\eta_{u}L_{f}}\ln\frac{2}{R^{2}}$ steps, we have $\|\bm{u}_{j}^{T}-\bm{v}_{i}^{T}\|_{2}\leq R$. This implies $\langle\bm{u}_{j}^{T},\bm{y}\rangle-\langle\bm{v}_{i}^{T},\bm{y}\rangle=\langle\bm{u}_{j}^{T}-\bm{v}_{i}^{T},\bm{y}\rangle\leq\|\bm{u}_{j}^{T}-\bm{v}_{i}^{T}\|\leq R$ and hence
$\Big{(}\langle\bm{v}_{i}^{T},\bm{y}\rangle-\langle\bm{v}_{i}^{0},\bm{y}\rangle\Big{)}-\Big{(}\langle\bm{u}_{j}^{T},\bm{y}\rangle-\langle\bm{u}_{j}^{0},\bm{y}\rangle\Big{)}=\Big{(}\langle\bm{u}_{j}^{0},\bm{y}\rangle-\langle\bm{v}_{i}^{0},\bm{y}\rangle\Big{)}-\Big{(}\langle\bm{u}_{j}^{T},\bm{y}\rangle-\langle\bm{v}_{i}^{T},\bm{y}\rangle\Big{)}\geq D-R.$ (10)
Multiplying (10) by $\frac{1+\eta_{u}}{\eta_{u}}$ and adding to (9):
$\Big{(}\frac{1+\eta_{c}}{\eta_{c}}+\frac{1+\eta_{u}}{\eta_{u}}\Big{)}\Big{(}\langle\bm{v}_{i}^{T},\bm{y}\rangle-\langle\bm{v}_{i}^{0},\bm{y}\rangle\Big{)}\geq\frac{1+\eta_{u}}{\eta_{u}}(D-R).$
This implies
$\langle\bm{v}_{i}^{T},\bm{y}\rangle-\langle\bm{v}_{i}^{0},\bm{y}\rangle\geq\frac{\frac{1+\eta_{u}}{\eta_{u}}}{\frac{1+\eta_{c}}{\eta_{c}}+\frac{1+\eta_{u}}{\eta_{u}}}(D-R)=\frac{\eta_{c}(1+\eta_{u})}{\eta_{u}(1+\eta_{c})+\eta_{c}(1+\eta_{u})}(D-R)\geq\frac{\eta_{c}}{\eta_{u}+\eta_{c}}(D-R).$
given $\eta_{c}\leq\eta_{u}$. ∎
The following lemma shows that, when we reflect some of the feature vectors in a system $(U^{t},V^{t})=(\{\bm{u}_{j}^{t}\}_{j\in[m]},\{\bm{v}_{i}^{t}\}_{i\in[n]})$, there is a correspondence between the behaviors of the system with the reflected vectors and the original system.
Lemma E.7 (Reflection).
Let $(U^{t},V^{t})=(\{\bm{u}_{j}^{t}\}_{j\in[m]},\{\bm{v}_{i}^{t}\}_{i\in[n]})$ be a system of $m$ users and $n$ creators with impact functions $f,g$. Let $a_{i},b_{j}\in\{+1,-1\}$, $\forall i\in[n]$, $\forall i\in[m]$ be some binary constants. Define:
$\tilde{\bm{u}}_{j}^{t}=b_{j}\bm{u}_{j}^{t}=\pm\bm{u}_{j}^{t},\quad\quad\tilde{\bm{v}}_{i}^{t}=a_{i}\bm{v}_{i}^{t}=\pm\tilde{\bm{v}}_{i}^{t}.$
and impact functions
$\tilde{f}(\tilde{\bm{v}}_{i},\tilde{\bm{u}}_{j})=a_{i}b_{j}f(\bm{v}_{i},\bm{u}_{j}),\quad\quad\tilde{g}(\tilde{\bm{u}}_{j},\tilde{\bm{v}}_{i})=a_{i}b_{j}g(\bm{u}_{j},\bm{v}_{i}).$
Then:
  • There is a “correspondence” between the evolution of the system $(U^{t},V^{t})$ with impact functions $f,g$ and the evolution of the system $(\tilde{U}^{t},\tilde{V}^{t})=(\{\tilde{\bm{u}}_{j}^{t}\}_{j\in[m]},\{\tilde{\bm{v}}_{i}^{t}\}_{i\in[n]})$ with impact functions $\tilde{f},\tilde{g}$. Formally, suppose every user is recommended the same creator in the two systems, then the updated vectors in the two systems still satisfy the relations: $\tilde{\bm{u}}_{j}^{t+1}=b_{j}\bm{u}_{j}^{t+1}$,  $\tilde{\bm{v}}_{i}^{t+1}=a_{i}\bm{v}_{i}^{t+1}$.
  • If the system $(\tilde{U}^{t},\tilde{V}^{t})$ is in $R$-bi-polarization, then the original system $(U^{t},V^{t})$ is also in $R$-bi-polarization.
Proof.
Consider the first item. Suppose user $i$ is recommended creator $j$ at time step $t$ in the two systems. Then by definition, the updated user vectors in the two systems satisfy
$\tilde{\bm{u}}_{j}^{t+1}$ $~{}=~{}\mathcal{P}\big{(}\tilde{\bm{u}}_{j}^{t}+\eta_{u}\tilde{f}(\tilde{\bm{v}}_{i}^{t},\tilde{\bm{u}}_{j}^{t})\tilde{\bm{v}}_{i}^{t}\big{)}~{}=~{}\mathcal{P}\big{(}b_{j}\bm{u}_{j}^{t}+\eta_{u}a_{i}b_{j}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})a_{i}\bm{v}_{i}^{t}\big{)}$
$~{}=~{}\mathcal{P}\big{(}b_{j}\bm{u}_{j}^{t}+\eta_{u}b_{j}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})\bm{v}_{i}^{t}\big{)}~{}=~{}b_{j}\mathcal{P}\big{(}\bm{u}_{j}^{t}+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})\bm{v}_{i}^{t}\big{)}~{}=~{}b_{j}\bm{u}_{j}^{t+1}$
Suppose creator $i$ is recommended to the set of users $J$ at time step $t$ in the two systems. Then,
$\tilde{\bm{v}}_{i}^{t+1}$ $~{}=~{}\mathcal{P}\big{(}\tilde{\bm{v}}_{i}^{t}+\tfrac{\eta_{c}}{|J|}\sum_{j\in J}g(\tilde{\bm{u}}_{j}^{t},\tilde{\bm{v}}_{i}^{t})\tilde{\bm{u}}_{j}^{t}\big{)}$
$~{}=~{}\mathcal{P}\big{(}a_{i}\bm{v}_{i}^{t}+\tfrac{\eta_{c}}{|J|}\sum_{j\in J}a_{i}b_{j}g(\bm{u}_{j}^{t},\bm{v}_{i}^{t})b_{j}\bm{u}_{j}^{t}\big{)}$
$~{}=~{}\mathcal{P}\big{(}a_{i}\bm{v}_{i}^{t}+\tfrac{\eta_{c}}{|J|}\sum_{j\in J}a_{i}g(\bm{u}_{j}^{t},\bm{v}_{i}^{t})\bm{u}_{j}^{t}\big{)}$
$~{}=~{}a_{i}\mathcal{P}\big{(}\bm{v}_{i}^{t}+\tfrac{\eta_{c}}{|J|}\sum_{j\in J}g(\bm{u}_{j}^{t},\bm{v}_{i}^{t})\bm{u}_{j}^{t}\big{)}~{}=~{}a_{i}\bm{v}_{i}^{t+1}.$
This means that the evolution of the system $(\tilde{U}^{t},\tilde{V}^{t})$ has a correspondence to the evolution of the original system $(U^{t},V^{t})$.
Consider the second item. Suppose $(\tilde{U}^{t},\tilde{V}^{t})$ is in $R$-bi-polarization, so $\tilde{\bm{v}}_{i}^{t}=\pm\bm{v}_{i}^{t}$ is $R$-close to ±$\pm\bm{c}$ and $\tilde{\bm{u}}_{j}^{t}=\pm\bm{u}_{j}^{t}$ is $R$-close to ±$\pm\bm{c}$ with some vector $\bm{c}\in\mathbb{S}^{d-1}$. This implies that $\bm{v}_{i}^{t}$ is $R$-close to ±$\pm\bm{c}$ and $\bm{u}_{j}^{t}$ is $R$-close to ±$\pm\bm{c}$. So, the system $(U^{t},V^{t})$ satisfies $R$-bi-polarization. ∎
Appendix F Proof of Lemma C.1
Lemma C.1 is proved by induction on the number $n$ of creators. We first show that any system with 1 creator and multiple users must converge to $R$-bi-polarization in finite steps for any $R>0$. Using the result for $1$ creator, we then construct a finite length path that leads to $R$-bi-polarization for any system with $n\geq 2$ creators.
F.1 Base Case: Convergence Results for $n=1$ Creator
We prove some convergence results for the special case of only one creator. This will serve as the basis for the proof for $n\geq 2$ creators. Recall that we have the following dynamics update rule:
  • User: $\bm{u}_{j}^{t+1}=\mathcal{P}(\bm{u}_{j}^{t}+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})\bm{v}_{i}^{t})$ where $\bm{v}_{i}^{t}$ is the creator recommended to user $j$; $f(\bm{v}_{i},\bm{u}_{j})$ satisfies:
    $f(\bm{v}_{i},\bm{u}_{j})\text{ is }\begin{cases}>0&\text{ if }\langle\bm{v}_{i},\bm{u}_{j}\rangle>0\\ <0&\text{ if }\langle\bm{v}_{i},\bm{u}_{j}\rangle<0\\ =0&\text{ if }\langle\bm{v}_{i},\bm{u}_{j}\rangle=0.\end{cases}$ (11)
  • Creator: $\bm{v}_{i}^{t+1}=\mathcal{P}(\bm{v}_{j}^{t}+\frac{\eta_{c}}{|J|}\sum_{j\in J}g(\bm{u}_{j}^{t},\bm{v}_{i}^{t})\bm{u}_{j}^{t})$ where $J$ is the set of users being recommended creator $i$.
Lemma F.1.
Consider a system of $1$ creator $\bm{v}_{i}^{t}$ and $|J|$ users $\{\bm{u}_{j}^{t}\}_{j\in J}$, where the creator is recommended to all users at every time step. Assume:
  • Initially, $\forall j\in J,\langle\bm{u}_{j}^{0},\bm{v}_{i}^{0}\rangle>0$.
  • There exists some constant $L_{f}>0$ such that $f(\bm{v}_{i},\bm{u}_{j})\geq L_{f}>0$ whenever $\langle\bm{v}_{i},\bm{u}_{j}\rangle>0$.
  • $g(\bm{u}_{j},\bm{v}_{i})=1$ when $\langle\bm{u}_{j},\bm{v}_{i}\rangle>0$.
  • $\eta_{c}\leq\frac{\eta_{u}L_{f}}{2}$ and $0\leq\eta_{u}<\frac{1}{2}$.
Then, for any $R>0$, after at most $\frac{8}{3\eta_{u}L_{f}}\ln\frac{2|J|}{R^{2}}$ steps, $\sum_{j\in J}\|\bm{u}_{j}^{t}-\bm{v}_{i}^{t}\|_{2}^{2}\leq R^{2}$ will hold forever. In particular, each user vector will satisfy $\|\bm{u}_{j}^{t}-\bm{v}_{i}^{t}\|_{2}\leq R$.
Proof.
We first note that, by Lemma E.5, all user vectors satisfy $\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle>0$ in all time steps $t>0$. Hence, the creator update is always $\bm{v}_{i}^{t+1}=\mathcal{P}(\bm{v}_{i}^{t}+\frac{\eta_{c}}{|J|}\sum_{j\in J}g(\bm{u}_{j}^{t},\bm{v}_{i}^{t})\bm{u}_{j}^{t})=\mathcal{P}(\bm{v}_{i}^{t}+\eta_{c}\frac{1}{|J|}\sum_{j\in J}\bm{u}_{j}^{t})$.
Let $a_{t}=1/(1-\frac{3\eta_{u}L_{f}}{8})^{t}$. Define the following potential function:
$\Phi^{t}~{}=~{}a_{t}\sum_{j\in J}\frac{1}{2}\|\bm{u}_{j}^{t}-\bm{v}_{i}^{t}\|_{2}^{2}~{}=~{}a_{t}\sum_{j\in J}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}.$ (12)
We will show that $\Phi^{t}$ is monotonically decreasing. Take the difference between $\Phi^{t+1}$ and $\Phi^{t}$:
$\Phi^{t+1}-\Phi^{t}~{}$ $=~{}a_{t+1}\sum_{j\in J}\Big{(}\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle-\langle\bm{u}_{j}^{t+1},\bm{v}_{i}^{t+1}\rangle\Big{)}~{}+~{}(a_{t+1}-a_{t})\sum_{j\in J}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}$
$=~{}a_{t+1}\bigg{(}\sum_{j\in J}\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}-\bm{u}_{j}^{t+1}\rangle+\sum_{j\in J}\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}-\bm{v}_{i}^{t+1}\rangle+\sum_{j\in J}\langle\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t},\bm{v}_{i}^{t}-\bm{v}_{i}^{t+1}\rangle\bigg{)}$
$\quad\quad+\,(a_{t+1}-a_{t})\sum_{j\in J}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}.$
Using Lemma E.3 with $\bm{x}^{t}=\bm{u}_{j}^{t}$, $\bm{z}^{t}=\bm{v}_{i}^{t}$, and $\eta=\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})$, we get
$\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}-\bm{u}_{j}^{t+1}\rangle~{}\leq~{}-\frac{\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})}{1+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}~{}\leq~{}-\frac{\eta_{u}L_{f}}{2}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}.$
Using Lemma E.4 with $\bm{x}^{t}=\bm{u}_{j}^{t}$, $\bm{z}^{t}=\bm{v}_{i}^{t}$, and $\eta=\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})$, we get
$\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}-\bm{u}_{j}^{t+1}\rangle~{}\leq~{}-\frac{1}{\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})}\|\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t}\|_{2}^{2}~{}\leq~{}-\frac{1}{\eta_{u}}\|\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t}\|_{2}^{2}.$
Using Lemma E.4 with $\bm{x}^{t}=\bm{v}_{i}^{t}$, $\bm{z}^{t}=\frac{1}{|J|}\sum_{j\in J}\bm{u}_{j}^{t}$, and $\eta=\eta_{c}$, we get
$\sum_{j\in J}\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}-\bm{v}_{i}^{t+1}\rangle~{}=~{}|J|\langle\tfrac{1}{|J|}\sum_{j\in J}\bm{u}_{j}^{t},\bm{v}_{i}^{t}-\bm{v}_{i}^{t+1}\rangle~{}\leq~{}-\frac{|J|}{\eta_{c}}\|\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t}\|_{2}^{2}.$
Using the above three inequalities, we can upper bound $\Phi^{t+1}-\Phi^{t}$:
$\Phi^{t+1}-\Phi^{t}$
$=~{}a_{t+1}\bigg{(}\frac{3}{4}\sum_{j\in J}\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}-\bm{u}_{j}^{t+1}\rangle~{}+~{}\frac{1}{4}\sum_{j\in J}\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}-\bm{u}_{j}^{t+1}\rangle$
$\hskip 40.00006pt\,+\,\sum_{j\in J}\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}-\bm{v}_{i}^{t+1}\rangle~{}+~{}\sum_{j\in J}\langle\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t},\bm{v}_{i}^{t}-\bm{v}_{i}^{t+1}\rangle\bigg{)}~{}+~{}(a_{t+1}-a_{t})\sum_{j\in J}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}$
$\leq~{}a_{t+1}\bigg{(}-\frac{3}{4}\sum_{j\in J}\frac{\eta_{u}L_{f}}{2}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}~{}-~{}\frac{1}{4}\sum_{j\in J}\frac{1}{\eta_{u}}\|\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t}\|_{2}^{2}$
$\hskip 40.00006pt\,-\,\frac{|J|}{\eta_{c}}\|\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t}\|_{2}^{2}~{}+~{}\sum_{j\in J}\|\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t}\|_{2}\cdot\|\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t}\|_{2}\bigg{)}\,+\,(a_{t+1}-a_{t})\sum_{j\in J}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}$
$=~{}a_{t+1}\bigg{(}-\frac{3\eta_{u}L_{f}}{8}\sum_{j\in J}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}$
$\hskip 40.00006pt~{}-~{}\sum_{j\in J}\Big{(}\underbrace{\frac{1}{4\eta_{u}}\|\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t}\|_{2}^{2}+\frac{1}{\eta_{c}}\|\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t}\|_{2}^{2}}_{\geq 2\sqrt{\frac{1}{4\eta_{u}\eta_{c}}\|\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t}\|_{2}^{2}\|\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t}\|_{2}^{2}}}~{}-~{}\|\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t}\|_{2}\cdot\|\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t}\|_{2}\Big{)}\bigg{)}$
$\quad\quad\,+\,(a_{t+1}-a_{t})\sum_{j\in J}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}$
$\leq~{}a_{t+1}\bigg{(}-\frac{3\eta_{u}L_{f}}{8}\sum_{j\in J}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}~{}-~{}\sum_{j\in J}\Big{(}\underbrace{\sqrt{\tfrac{1}{\eta_{u}\eta_{c}}}-1}_{\geq 0}\Big{)}\|\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t}\|_{2}\cdot\|\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t}\|_{2}\bigg{)}$
$\quad\quad\,+\,(a_{t+1}-a_{t})\sum_{j\in J}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}$
$\leq~{}a_{t+1}\bigg{(}-\frac{3\eta_{u}L_{f}}{8}\sum_{j\in J}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}~{}+~{}0\bigg{)}~{}+~{}(a_{t+1}-a_{t})\sum_{j\in J}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}$
$=~{}\Big{(}\big{(}1-\frac{3\eta_{u}L_{f}}{8}\big{)}a_{t+1}-a_{t}\Big{)}\sum_{j\in J}\big{(}1-\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle\big{)}$
$=~{}0$
where the last step is because $(1-\frac{3\eta_{u}L_{f}}{8})a_{t+1}=a_{t}$.
We have shown that $\Phi^{t}$ is monotonically decreasing. Thus,
$\frac{1}{2}\sum_{j\in J}\|\bm{u}_{j}^{T}-\bm{v}_{i}^{T}\|^{2}\,=\,\frac{\Phi^{T}}{a_{T}}\,\leq\,\frac{\Phi^{0}}{a_{T}}\,\leq\,\frac{\sum_{j\in J}1}{a_{T}}\,=\,\big{(}1-\tfrac{3\eta_{u}L_{f}}{8}\big{)}^{T}|J|\,\leq\,e^{-\frac{3\eta_{u}L_{f}}{8}T}|J|\,\leq\,\frac{1}{2}R^{2}$
whenever $T\geq\frac{8}{3\eta_{u}L_{f}}\ln\frac{2|J|}{R^{2}}$. ∎
Corollary F.2 (of Lemma F.1).
Consider a system of $1$ creator $\bm{v}_{i}^{t}$ and $|J|$ users $\{\bm{u}_{j}^{t}\}_{j\in J}$, where the creator is recommended to all users at every time step. Assume:
  • Initially, $\langle\bm{u}_{j}^{0},\bm{v}_{i}^{0}\rangle\neq 0$ for every $j\in J$.
  • There exists some constant $L_{f}>0$ such that $|f(\bm{v}_{i},\bm{u}_{j})|\geq L_{f}>0$ whenever $\langle\bm{v}_{i},\bm{u}_{j}\rangle\neq 0$.
  • $g(\bm{u}_{j},\bm{v}_{i})=\mathrm{sign}(\langle\bm{u}_{j},\bm{v}_{i}\rangle)$.
  • $\eta_{c}\leq\frac{\eta_{u}L_{f}}{2}$ and $0\leq\eta_{u}<\frac{1}{2}$.
Then, for any $R>0$, after at most $\frac{8}{3\eta_{u}L_{f}}\ln\frac{2|J|}{R^{2}}$ steps, the system will reach $R$-bi-polarization.
Proof.
Let $J^{+}=\{j\in J:\langle\bm{u}_{j}^{0},\bm{v}_{i}^{0}\rangle>0\}$ be the set of users with positive inner products with creator $i$ initially; let $J^{-}=\{j\in J:\langle\bm{u}_{j}^{0},\bm{v}_{i}^{0}\rangle<0\}$. Let $\tilde{\bm{u}}_{j}^{t}=-\bm{u}_{j}^{t}$ for $j\in J^{-}$ and $\tilde{\bm{u}}_{j}^{t}=\bm{u}_{j}^{t}$ for $j\in J^{+}$. Then, the system consisting of $\{\tilde{\bm{u}}_{j}^{t}\}_{j\in J}$ and $\bm{v}_{i}^{t}$ satisfies the initial condition $\langle\tilde{\bm{u}}_{j}^{0},\bm{v}_{i}^{0}\rangle>0$ in Lemma F.1. So, by Lemma F.1, it reaches $R$-consensus after at most $\frac{8}{3\eta_{u}L_{f}}\ln\frac{2|J|}{R^{2}}$ steps. Then by the reflection lemma (Lemma E.7), the original system, consisting of $\{\bm{u}_{j}^{t}\}_{j\in J}$ and $\bm{v}_{i}^{t}$, must reach $R$-bi-polarization. ∎
F.2 Inductive Step: Proof of Lemma C.1
Lemma F.3.
Consider a system of $n\geq 1$ creators $\{\bm{v}_{1}^{t},\ldots,\bm{v}_{n}^{t}$} and $|J|$ users $\{\bm{u}_{j}^{t}\}_{j\in J}$. Assume:
  • Initially, $\langle\bm{v}_{i}^{0},\bm{v}_{i^{\prime}}^{0}\rangle>0$ for every $i,i^{\prime}$, and $\langle\bm{v}_{i}^{0},\bm{u}_{j}^{0}\rangle>0$ for every $i,j$.
  • Assumptions of Lemma F.1.
Then, for any $R\in(0,1)$, there exists a path of finite length that leads the initial state $(\bm{U}^{0},\bm{V}^{0})$ to $R$-consensus.
Proof.
Fix any $R\in(0,1)$. Choose $R_{1}$ such that $\sqrt{(\frac{\eta_{u}}{\eta_{c}}+2)4R_{1}}=R$. Clearly, $R_{1}<R$. We construct a path that leads the state $(\bm{U}^{0},\bm{V}^{0})$ to $R$-consensus as follows.
Step (1): Consider the subsystem of the first $n-1$ creators and all users $J$. By induction, there exists a path of length $T_{1}=L_{n-1,R_{1}}<+\infty$ that leads the subsystem to $(R_{1},\bm{c}^{T_{1}})$-consensus with some $\bm{c}^{T_{1}}\in\mathbb{S}^{d-1}$. So, after these $T_{1}$ steps, all creators $i\in\{1,\ldots,n-1\}$ and all users $j\in J$ satisfy $\|\bm{v}_{i}^{T_{1}}-\bm{c}^{T_{1}}\|\leq R_{1}$ and $\|\bm{u}_{j}^{T_{1}}-\bm{c}^{T_{1}}\|\leq R_{1}$. Creator $n$ does not update during these $T_{1}$ steps, so $\bm{v}_{n}^{T_{1}}=\bm{v}_{n}^{0}$, and it still has positive inner products with the first $n-1$ creators and all users by the convex cone property (Lemma E.2). Let’s then consider the distance between creators $n$ and the consensus center $\bm{c}^{T_{1}}$: $\|\bm{v}_{n}^{T_{1}}-\bm{c}^{T_{1}}\|$. If $\|\bm{v}_{n}^{T_{1}}-\bm{c}^{T_{1}}\|\leq R$, then the system has satisfied $(R,\bm{c}^{T_{1}})$-consensus, so our construction is finished. Otherwise, $\|\bm{v}_{n}^{T_{1}}-\bm{c}^{T_{1}}\|>R$. We continue the construction as follows:
Step (2): Pick any user $j_{0}\in J$, recommend creator $n$ to user $j_{0}$ for $T_{2}=\frac{8}{3\eta_{u}L_{f}}\ln\frac{2}{R_{1}^{2}}$ steps, while recommending creator 1 to all other users. From the $(R_{1},\bm{c}^{T_{1}})$-consensus in step (1) we know $\|\bm{u}_{j_{0}}^{T_{1}}-\bm{c}^{T_{1}}\|\leq R_{1}$, so
$\langle\bm{u}_{j_{0}}^{T_{1}},\bm{c}^{T_{1}}\rangle~{}=~{}1-\tfrac{1}{2}\|\bm{u}_{j_{0}}^{T_{1}}-\bm{c}^{T_{1}}\|^{2}~{}\geq~{}1-\tfrac{R_{1}^{2}}{2}~{}>~{}1-\tfrac{R^{2}}{2}~{}\geq~{}1-\tfrac{1}{2}\|\bm{v}_{n}^{T_{1}}-\bm{c}^{T_{1}}\|^{2}~{}=~{}\langle\bm{v}_{2}^{T_{1}},\bm{c}^{T_{1}}\rangle.$
Thus, we can apply Lemma E.6 with $\bm{y}=\bm{c}^{T_{1}}$ to derive that, after these $T_{2}$ steps,
$\langle\bm{v}_{n}^{T_{1}+T_{2}},\bm{c}^{T_{1}}\rangle-\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle$ $~{}\geq~{}\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}\langle\bm{u}_{j_{0}}^{T_{1}},\bm{c}^{T_{1}}\rangle-\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle-R_{1}\Big{)}$
$~{}\geq~{}\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}1-\tfrac{R_{1}^{2}}{2}-\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle-R_{1}\Big{)}.$
$\implies\quad\langle\bm{v}_{n}^{T_{1}+T_{2}},\bm{c}^{T_{1}}\rangle$ $~{}\geq~{}\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}1-\tfrac{R_{1}^{2}}{2}-\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle-R_{1}\Big{)}.$ (13)
For the inner product between creator $n$ and user $j_{0}$, by Lemma F.1 $\|\bm{v}_{n}^{T_{1}+T_{2}}-\bm{u}_{j_{0}}^{T_{1}+T_{2}}\|\leq R_{1}$, so
$\langle\bm{v}_{n}^{T_{1}+T_{2}},\bm{u}_{j_{0}}^{T_{1}+T_{2}}\rangle~{}=~{}1-\tfrac{1}{2}\|\bm{v}_{n}^{T_{1}+T_{2}}-\bm{u}_{j_{0}}^{T_{1}+T_{2}}\|^{2}~{}\geq~{}1-\tfrac{R_{1}^{2}}{2}.$ (14)
Consider the inner products between creator $n$ and the first $n-1$ creators and the users in $J\setminus\{j_{0}\}$. Because the first $n-1$ creators and the users in $J\setminus\{j_{0}\}$ form $(R_{1},\bm{c}^{T_{1}})$-consensus at time step $T_{1}$, by Observation 3.2, they still form $(R_{1},\bm{c}^{T_{1}})$-consensus at time step $T_{1}+T_{2}$, so $\|\bm{v}_{i}^{T_{1}+T_{2}}-\bm{c}^{T_{1}}\|\leq R_{1}$ and $\|\bm{u}_{j}^{T_{1}+T_{2}}-\bm{c}^{T_{1}}\|\leq R_{1}$. This implies, for $i\neq n$,
$\langle\bm{v}_{n}^{T_{1}+T_{2}},\bm{v}_{i}^{T_{1}+T_{2}}\rangle$ $~{}\geq~{}\langle\bm{v}_{n}^{T_{1}+T_{2}},\bm{c}^{T_{1}}\rangle-\|\bm{v}_{i}^{T_{1}+T_{2}}-\bm{c}^{T_{1}}\|$
$~{}\geq~{}\langle\bm{v}_{2}^{T_{1}+T_{2}},\bm{c}^{T_{1}}\rangle-R_{1}$
(13) $~{}\geq~{}\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}1-\tfrac{R_{1}^{2}}{2}-\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle-R_{1}\Big{)}-R_{1},$ (15)
and for $j\in J\setminus\{j_{0}\}$,
$\langle\bm{v}_{n}^{T_{1}+T_{2}},\bm{u}_{j}^{T_{1}+T_{2}}\rangle$ $~{}\geq~{}\langle\bm{v}_{n}^{T_{1}+T_{2}},\bm{c}^{T_{1}}\rangle-\|\bm{u}_{j}^{T_{1}+T_{2}}-\bm{c}^{T_{1}}\|$
$~{}\geq~{}\langle\bm{v}_{2}^{T_{1}+T_{2}},\bm{c}^{T_{1}}\rangle-R_{1}$
(13) $~{}\geq~{}\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}1-\tfrac{R_{1}^{2}}{2}-\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle-R_{1}\Big{)}-R_{1}.$ (16)
Step (3): Consider the subsystem of the first $n-1$ creators and all users $J$. By induction, there exists a path of length $T_{3}=L_{n-1,R_{1}}<+\infty$ that leads the subsystem to $(R_{1},\bm{c}^{T_{1}+T_{2}+T_{3}})$-consensus with some $\bm{c}^{T_{1}+T_{2}+T_{3}}\in\mathbb{S}^{d-1}$. So, we have $\|\bm{v}_{i}^{T_{1}+T_{2}+T_{3}}-\bm{c}^{T_{1}+T_{2}+T_{3}}\|\leq R_{1}$ for every $i\in\{1,\ldots,n-1\}$ and $\|\bm{u}_{j}^{T_{1}+T_{2}+T_{3}}-\bm{c}^{T_{1}+T_{2}+T_{3}}\|\leq R_{1}$ for every $j\in J$, and $\bm{v}_{n}^{T_{1}+T_{2}+T_{3}}=\bm{v}_{n}^{T_{1}+T_{2}}$. Consider the inner product between creator $n$ and any of the first $n-1$ creators $i\in\{1,\ldots,n-1\}$. By the convex cone property (Lemma E.2),
$\langle\bm{v}_{n}^{T_{1}+T_{2}+T_{3}},\bm{v}_{i}^{T_{1}+T_{2}+T_{3}}\rangle$ $~{}=~{}\langle\bm{v}_{n}^{T_{1}+T_{2}},\bm{v}_{i}^{T_{1}+T_{2}+T_{3}}\rangle$
Lemma E.2 $~{}\geq~{}\min\big{\{}\langle\bm{v}_{n}^{T_{1}+T_{2}},\bm{v}_{i}^{T_{1}+T_{2}}\rangle,~{}\min_{j\in J}\langle\bm{v}_{n}^{T_{1}+T_{2}},\bm{u}_{j}^{T_{1}+T_{2}}\rangle\big{\}}$
by (14), (15), (16) $~{}\geq~{}\min\big{\{}\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}1-\tfrac{R_{1}^{2}}{2}-\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle-R_{1}\Big{)}-R_{1},~{}1-\tfrac{R_{1}^{2}}{2}\big{\}}$
$~{}=~{}\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}1-\tfrac{R_{1}^{2}}{2}-\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle-R_{1}\Big{)}-R_{1}$ (17)
where the last equality is because, under the assumption of $\|\bm{v}_{n}^{T_{1}}-\bm{c}^{T_{1}}\|>R=\sqrt{(\frac{\eta_{u}}{\eta_{c}}+2)4R_{1}}$,
$\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}1-\tfrac{R_{1}^{2}}{2}-\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle-R_{1}\Big{)}-R_{1}$
$~{}=~{}\tfrac{\eta_{u}}{\eta_{u}+\eta_{c}}\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}1-\tfrac{R_{1}^{2}}{2}-R_{1}\Big{)}-R_{1}$
$~{}=~{}\tfrac{\eta_{u}}{\eta_{u}+\eta_{c}}\Big{(}1-\tfrac{1}{2}\|\bm{v}_{2}^{T_{1}}-\bm{c}^{T_{1}}\|^{2}\Big{)}+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}1-\tfrac{R_{1}^{2}}{2}-R_{1}\Big{)}-R_{1}$
$~{}\leq~{}\tfrac{\eta_{u}}{\eta_{u}+\eta_{c}}\Big{(}1-\tfrac{1}{2}(\tfrac{\eta_{u}}{\eta_{c}}+2)4R_{1}\Big{)}+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}1-\tfrac{R_{1}^{2}}{2}-R_{1}\Big{)}-R_{1}$
$~{}\leq~{}\max\{1-\tfrac{1}{2}(\tfrac{\eta_{u}}{\eta_{c}}+2)4R_{1},~{}1-\tfrac{R_{1}^{2}}{2}-R_{1}\}-R_{1}$
$~{}=~{}1-\tfrac{R_{1}^{2}}{2}-R_{1}-R_{1}~{}\leq~{}1-\tfrac{R_{1}^{2}}{2}.$
From (17) and $\|\bm{v}_{i}^{T_{1}+T_{2}+T_{3}}-\bm{c}^{T_{1}+T_{2}+T_{3}}\|\leq R_{1}$,
$\langle\bm{v}_{n}^{T_{1}+T_{2}+T_{3}},\bm{c}^{T_{1}+T_{2}+T_{3}}\rangle$ $~{}\geq~{}\langle\bm{v}_{n}^{T_{1}+T_{2}},\bm{v}_{i}^{T_{1}+T_{2}+T_{3}}\rangle-\|\bm{v}_{i}^{T_{1}+T_{2}+T_{3}}-\bm{c}^{T_{1}+T_{2}+T_{3}}\|$
$~{}\geq~{}\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}1-\tfrac{R_{1}^{2}}{2}-\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle-R_{1}\Big{)}-2R_{1}.$
Using $1$ to minus the above inequality, we obtain
$1-\langle\bm{v}_{n}^{T_{1}+T_{2}+T_{3}},\bm{c}^{T_{1}+T_{2}+T_{3}}\rangle$ $~{}\leq~{}\tfrac{\eta_{u}}{\eta_{u}+\eta_{c}}\Big{(}1-\langle\bm{v}_{n}^{T_{1}},\bm{c}^{T_{1}}\rangle\Big{)}+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}\tfrac{R_{1}^{2}}{2}+R_{1}\Big{)}+2R_{1}.$
Let $F^{t}=1-\langle\bm{v}_{n}^{t},\bm{c}^{t}\rangle$, then
$F^{T_{1}+T_{2}+T_{3}}$ $\leq\tfrac{\eta_{u}}{\eta_{u}+\eta_{c}}F^{T_{1}}+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}\tfrac{R_{1}^{2}}{2}+R_{1}\Big{)}+2R_{1}.$ (18)
Repeat steps (2) and (3) for $K$ times. Then, using (18) for $K$ times,
$F^{T_{1}+K(T_{2}+T_{3})}$ $~{}\leq~{}\tfrac{\eta_{u}}{\eta_{u}+\eta_{c}}F^{T_{1}+(K-1)(T_{2}+T_{3})}+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}\tfrac{R_{1}^{2}}{2}+R_{1}\Big{)}+2R_{1}$
$~{}\leq~{}\tfrac{\eta_{u}}{\eta_{u}+\eta_{c}}\bigg{(}\tfrac{\eta_{u}}{\eta_{u}+\eta_{c}}F^{T_{1}+(K-2)(T_{2}+T_{3})}+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}\tfrac{R_{1}^{2}}{2}+R_{1}\Big{)}+2R_{1}\bigg{)}+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}\tfrac{R_{1}^{2}}{2}+R_{1}\Big{)}+2R_{1}$
[…]
$~{}\leq~{}\big{(}\tfrac{\eta_{u}}{\eta_{u}+\eta_{c}}\big{)}^{K}F^{T_{1}}+\Big{(}1+\tfrac{\eta_{u}}{\eta_{t}+\eta_{c}}+\cdots+\big{(}\tfrac{\eta_{u}}{\eta_{t}+\eta_{c}}\big{)}^{K-1}\Big{)}\Big{(}\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}\tfrac{R_{1}^{2}}{2}+R_{1}\Big{)}+2R_{1}\Big{)}$
$~{}\leq~{}\big{(}\tfrac{\eta_{u}}{\eta_{u}+\eta_{c}}\big{)}^{K}\cdot 1+\frac{1}{1-\tfrac{\eta_{u}}{\eta_{u}+\eta_{c}}}\Big{(}\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}\tfrac{R_{1}^{2}}{2}+R_{1}\Big{)}+2R_{1}\Big{)}$
$~{}=~{}\big{(}\tfrac{\eta_{u}}{\eta_{u}+\eta_{c}}\big{)}^{K}+\tfrac{R_{1}^{2}}{2}+R_{1}+\tfrac{\eta_{u}+\eta_{c}}{\eta_{c}}2R_{1}$
$~{}\leq~{}\tfrac{R_{1}^{2}}{2}+\tfrac{R_{1}^{2}}{2}+R_{1}+\tfrac{\eta_{u}+\eta_{c}}{\eta_{c}}2R_{1}~{}\leq~{}\big{(}\tfrac{\eta_{u}}{\eta_{c}}+2)2R_{1},$
by choosing $K=\frac{\ln\frac{2}{R_{1}^{2}}}{\ln\frac{\eta_{u}+\eta_{c}}{\eta_{u}}}\leq\tfrac{\eta_{u}+\eta_{c}}{\eta_{c}}\ln\tfrac{2}{R_{1}^{2}}$. This means that, after repeating steps (2) and (3) for $K$ times, we must have
$\|\bm{v}_{n}^{T_{1}+K(T_{2}+T_{3})}-\bm{c}^{T_{1}+K(T_{2}+T_{3})}\|$ $~{}=~{}\sqrt{2\big{(}1-\langle\bm{v}_{n}^{T_{1}+K(T_{2}+T_{3})},\bm{c}^{T_{1}+K(T_{2}+T_{3})}\rangle\big{)}}$
$~{}=~{}\sqrt{2F^{T_{1}+K(T_{2}+T_{3})}}~{}\leq~{}\sqrt{2\big{(}\tfrac{\eta_{u}}{\eta_{c}}+2)2R_{1}}~{}=~{}R.$
The above inequality, together with the fact that other creators $i\neq n$ and all users in $J$ already satisfy $(R_{1}\leq R,\bm{c}^{T_{1}+K(T_{2}+T_{3})})$-consensus after step (3), implies that the whole system has reached $(R,\bm{c}^{T_{1}+K(T_{2}+T_{3})})$-consensus.
The length of the path constructed above is at most:
$T_{1}+K(T_{2}+T_{3})$ $~{}\leq~{}L_{n-1,R_{1}}+\tfrac{\eta_{u}+\eta_{c}}{\eta_{c}}\ln\tfrac{2}{R_{1}^{2}}\Big{(}\tfrac{8}{3\eta_{u}L_{f}}\ln\tfrac{2|J|}{R_{1}^{2}}+L_{n-1,R_{1}}\Big{)}~{}=~{}L_{n,R}~{}<~{}+\infty,$
which is finite. ∎
Lemma F.4.
Consider a subsystem of $n$ creators $\{\bm{v}_{1}^{t},\ldots,\bm{v}_{n}^{t}$} and $|J|$ users $\{\bm{u}_{j}^{t}\}_{j\in J}$. Assume:
  • Initially, the first $n-1$ creators and all users are in $R_{0}$-consensus: $\|\bm{v}_{i}^{0}-\bm{c}\|\leq R_{0}$, $\|\bm{u}_{j}^{0}-\bm{c}\|\leq R_{0}$, with $0<R_{0}<\frac{\eta_{c}}{5(\eta_{c}+\eta_{u})}$.
  • $\langle\bm{v}_{n}^{0},\bm{u}_{j_{0}}^{0}\rangle>0$ for some $j_{0}\in J$.
  • $g(\bm{u}_{j},\bm{v}_{i})=\mathrm{sign}(\langle\bm{u}_{j},\bm{v}_{i}\rangle)$.
  • Assumption of Lemma F.1.
Then, for any $R\in(0,1)$, there exists a path of finite length that leads the initial state $(\bm{U}^{0},\bm{V}^{0})$ to $R$-consensus.
Proof.
First, we recommend creator $n$ to user $j_{0}$ for $T=\frac{8}{3\eta_{u}L_{f}}\ln\frac{2}{R_{0}^{2}}$ steps, while recommending other creators to other users arbitrarily. Applying Lemma E.6 with $\bm{y}=\bm{u}_{j_{0}}^{0}$, we get
$\langle\bm{v}_{n}^{T},\bm{u}_{j_{0}}^{0}\rangle-\langle\bm{v}_{n}^{0},\bm{u}_{j_{0}}^{0}\rangle~{}\geq~{}\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}\langle\bm{u}_{j_{0}}^{0},\bm{u}_{j_{0}}^{0}\rangle-\langle\bm{v}_{n}^{0},\bm{u}_{j_{0}}^{0}\rangle-R_{0}\Big{)}~{}=~{}\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}1-\langle\bm{v}_{n}^{0},\bm{u}_{j_{0}}^{0}\rangle-R_{0}\Big{)}.$ (19)
On the other hand, because the first $n-1$ creators and all users in $J\setminus\{j_{0}\}$ form an $(R_{0},\bm{c})$-consensus at time step 00, according to Observation 3.2, they still form an $(R_{0},\bm{c})$-consensus at time step $T$, so $\|\bm{v}_{i}^{T}-\bm{c}\|\leq R_{0}$ for every $i\in\{1,\ldots,n-1\}$. This implies, for every $i\in\{1,\ldots,n-1\}$,
$\langle\bm{v}_{n}^{T},\bm{v}_{i}^{T}\rangle-\langle\bm{v}_{n}^{T},\bm{u}_{j_{0}}^{0}\rangle~{}\geq~{}-\|\bm{v}_{i}^{T}-\bm{u}_{j_{0}}^{0}\|~{}\geq~{}-\|\bm{v}_{i}^{T}-\bm{c}\|-\|\bm{c}-\bm{u}_{j_{0}}^{0}\|~{}\geq~{}-2R_{0}.$ (20)
Adding (19) and (20) and moving $\langle\bm{v}_{n}^{0},\bm{u}_{j_{0}}^{0}\rangle$ to the right side, we get
$\langle\bm{v}_{n}^{T},\bm{v}_{i}^{T}\rangle$ $~{}\geq~{}\langle\bm{v}_{n}^{0},\bm{u}_{j_{0}}^{0}\rangle+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\Big{(}1-\langle\bm{v}_{n}^{0},\bm{u}_{j_{0}}^{0}\rangle-R_{0}\Big{)}-2R_{0}$
$~{}=~{}\tfrac{\eta_{u}}{\eta_{u}+\eta_{c}}\langle\bm{v}_{n}^{0},\bm{u}_{j_{0}}^{0}\rangle+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\big{(}1-R_{0}\big{)}-2R_{0}$
$~{}>~{}0+\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\big{(}1-R_{0}\big{)}-2R_{0}~{}>~{}0,$
under the condition of $R_{0}<\frac{\eta_{c}}{5(\eta_{u}+\eta_{c})}$. Moreover, for every $j\in J\setminus\{j_{0}\}$, because $\|\bm{u}_{j}^{T}-\bm{v}_{i}^{T}\|\leq\|\bm{u}_{j}^{T}-\bm{c}\|+\|\bm{c}-\bm{v}_{i}^{T}\|\leq 2R_{0}$,
$\langle\bm{v}_{n}^{T},\bm{u}_{j}^{T}\rangle~{}\geq~{}\langle\bm{v}_{n}^{T},\bm{v}_{i}^{T}\rangle-\|\bm{u}_{j}^{T}-\bm{v}_{i}^{T}\|~{}\geq~{}\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\big{(}1-R_{0}\big{)}-4R_{0}>0.$
For $j_{0}$, by Lemma F.1, $\|\bm{v}_{n}^{T}-\bm{u}_{j_{0}}^{T}\|\leq R_{0}$, so
$\langle\bm{v}_{n}^{T},\bm{u}_{j_{0}}^{T}\rangle~{}=~{}1-\tfrac{1}{2}\|\bm{v}_{n}^{T}-\bm{u}_{j_{0}}^{T}\|^{2}~{}\geq~{}1-\tfrac{R_{0}^{2}}{2}~{}>~{}0.$
For the inner product between any creator $i\in\{1,\ldots,n-1\}$ and the users:
$\langle\bm{v}_{i}^{T},\bm{u}_{j_{0}}^{T}\rangle~{}\geq~{}\langle\bm{v}_{i}^{T},\bm{v}_{n}^{T}\rangle-\|\bm{v}_{n}^{T}-\bm{u}_{j_{0}}^{T}\|~{}\geq~{}\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\big{(}1-R_{0}\big{)}-2R_{0}-R_{0}~{}=~{}\tfrac{\eta_{c}}{\eta_{u}+\eta_{c}}\big{(}1-R_{0}\big{)}-3R_{0}~{}>~{}0;$
$\forall j\in J\setminus\{j_{0}\},\quad\langle\bm{v}_{i}^{T},\bm{u}_{j}^{T}\rangle~{}=~{}1-\tfrac{1}{2}\|\bm{v}_{i}^{T}-\bm{u}_{j}^{T}\|^{2}~{}\geq~{}1-\tfrac{1}{2}\big{(}\|\bm{v}_{i}^{T}-\bm{c}\|+\|\bm{c}-\bm{u}_{j}^{T}\|\big{)}^{2}~{}>~{}1-\tfrac{1}{2}(2R_{0})^{2}~{}>~{}0.$
All of the “$>0$” inequalities above show that the system of $\{\bm{v}_{i}^{T}\}_{i\in[n]}$ and $\{\bm{u}_{j}^{T}\}_{j\in J}$ satisfies the condition of Lemma F.3. So, there exists a path of finite length $T_{2}<+\infty$ that leads the system to $R$-consensus by Lemma F.3. The total length of path $T+T_{2}=\tfrac{8}{3\eta_{u}L_{f}}\ln\tfrac{2}{R_{0}^{2}}+T_{2}<+\infty$ is finite. ∎
See C.1
Proof.
We prove this lemma by induction on the number of creators $n$. The case for $n=1$ directly follows from Corollary F.2 which shows that, for any system of $n=1$ creator and $|J|$ users with no $\langle\bm{v}_{i}^{0},\bm{u}_{j}^{0}\rangle=0$, there exists a path of length at most $L_{1}^{R}=\frac{8}{3\eta_{u}L_{F}}\ln\frac{2|J|}{R^{2}}<+\infty$ that leads to $R$-bi-polarization.
Consider $n\geq 2$. Consider the subsystem consisting of the first $n-1$ creators $\{\bm{v}_{1}^{t},\ldots,\bm{v}_{n-1}^{t}\}$ and all users. Let $R_{0}=\frac{\eta_{c}}{6(\eta_{c}+\eta_{u})}$. By induction, there exists a path of finite length $T_{1}=L_{n-1}^{R_{0}}<+\infty$ that leads the subsystem to $R_{0}$-bi-polarization, with some vector $\bm{c}_{0}\in\mathbb{S}^{d-1}$, so every $\bm{v}_{i}^{T_{1}}$ is $R_{0}$-close to $+\bm{c}_{0}$ or $-\bm{c}_{0}$, for $i\neq n$, and every $\bm{u}_{j}^{T_{1}}$ is $R_{0}$-close to $+\bm{c}_{0}$ or $-\bm{c}_{0}$. Define:
$\tilde{\bm{v}}_{i}^{t}=\begin{cases}\bm{v}_{i}^{t}&\text{ if $\bm{v}_{i}^{T_{1}}$ is $R_{0}$-close to $+\bm{c}$}\\ -\bm{v}_{i}^{t}&\text{ if $\bm{v}_{i}^{T_{1}}$ is $R_{0}$-close to $-\bm{c}$}\end{cases}~{}~{}~{}~{}\forall i\neq n,\quad\quad\quad\tilde{\bm{u}}_{j}^{t}=\begin{cases}\bm{u}_{j}^{t}&\text{ if $\bm{u}_{j}^{T_{1}}$ is $R_{0}$-close to$+\bm{c}$}\\ -\bm{u}_{j}^{t}&\text{ if $\bm{u}_{j}^{T_{1}}$ is $R_{0}$-close to $-\bm{c}$}\end{cases}~{}~{}~{}~{}\forall j\in J.$
By definition, we have
$\|\tilde{\bm{v}}_{i}^{T_{1}}-\bm{c}_{0}\|\leq R_{0},~{}~{}~{}\forall i\neq n,~{}\quad\quad\quad\|\tilde{\bm{u}}_{j}^{T_{1}}-\bm{c}_{0}\|\leq R_{0},~{}~{}~{}\forall j\in J.$
This means that $\{\tilde{\bm{v}}_{i}^{T_{1}}\}_{i\neq n}$ and $\{\tilde{\bm{u}}_{j}^{T_{1}}\}_{j\in J}$ form an $(R_{0},\bm{c}_{0})$-consensus. Consider creator $n$. Let
$\tilde{\bm{v}}_{n}^{t}=\begin{cases}\bm{v}_{n}^{t}&\text{ if $\langle\bm{v}_{n}^{T_{1}},\tilde{\bm{u}}_{j_{0}}^{T_{1}}\rangle>0$ for some $j_{0}\in J$}\\ -\bm{v}_{n}^{t}&\text{ if $\langle\bm{v}_{n}^{T_{1}},\tilde{\bm{u}}_{j}^{T_{1}}\rangle<0$ for all $j\in J$}.\end{cases}$
(The case where $\langle\bm{v}_{n}^{T_{1}},\tilde{\bm{u}}_{j}^{T_{1}}\rangle=0$ for some $j\in J$ is ignored because the initial states that can lead to such states have measure 00.) By definition, we have
$\langle\tilde{\bm{v}}_{n}^{T_{1}},\tilde{\bm{u}}_{j_{0}}^{T_{1}}\rangle>0\text{ for some $j_{0}\in J$}.$
Note that, at time step $T_{1}$, the system consisting of $\{\tilde{\bm{v}}_{i}^{T_{1}}\}_{i\in[n]}$ and $\{\tilde{\bm{u}}_{j}^{T_{1}}\}_{j\in J}$ satisfies the condition of Lemma F.4, so there exists a path of length $T_{2}=\tilde{L}_{n}^{R}<+\infty$ that leads the system to $R$-consensus. Then by the reflection lemma (Lemma E.7), the original system $\{\bm{v}_{i}^{t}\}_{i\in[n]}$, $\{\bm{u}_{j}^{t}\}_{j\in J}$ must reach $R$-bi-polarization. The total length of path that leads to this $R$-bi-polarization is $L_{n}^{R}=T_{1}+T_{2}=L_{n-1}^{R_{0}}+\tilde{L}_{n}^{R}<+\infty$. ∎
Appendix G Additional Discussion on Real-World Recommender Systems
Here we further discuss real-world recommender systems’ properties and designs that are currently not covered in our main paper. We plan to generalize our model in the future to further capture these features and discuss insightful findings, but having them in the current paper may be a distraction to our main findings.
G.1 User and Creator Retention and Activeness
In our current model, the users and creators will stay in the system from the start to the end. However, in real-world recommender systems, users and creators may leave the platform either permanently or for a certain period. Meanwhile, new users and creators will join the platform. Such join and leave dynamics are also influenced by the recommendations’ relevance and diversity, which further complicate the problem. Moreover, users and creators have different activeness levels on the platform, e.g., some users may watch a lot more videos than others, and some creators may post a lot more creations, these effects will also be strongly correlated with the dual influence of the recommender system.
G.2 Creation Quality
Creation quality is a major factor influencing users’ feedback in addition to the creation style, e.g., well-made cuisine videos could also be fun and liked by gamers and pet lovers, which we need more than a collaborative filtering type of modeling like our current model to capture such features. A potential solution to boost both long-term system diversity and single-shot recommendation diversity is to design mechanisms that can incentivize creators to create higher-quality videos instead of changing their creation styles.
G.3 Cold Start
Cold Start is widely used in real-world recommender systems for newly published items. Due to the lack of user-item interactions on new items, the systems randomly recommend these new items to users and collect data for collaborative filtering. In our current model, if we consider the creators creating new items in each time step under their current time creation style, then cold start guarantees the conditions in Thm 3.3. But if we consider the system to have good enough content understanding ability and can accurately predict the new creations’ embeddings, the cold start is not necessary and our model and results in the top-$k$ recommendation and truncation parts are valid. We also highlight a subtle difference between cold start and random traffic, if cold start is used on creators instead of items, then after the creator is exposed to users a certain number of times, the system will not guarantee to provide a non-zero probability of recommending this creator, and thus the conditions in Thm 3.3 may not hold.